T.R | Title | User | Personal Name | Date | Lines |
---|
45.52 | brush height problem; lift arm the culprit | SUBPAC::LANDRY | | Sun Nov 12 1995 19:17 | 13 |
| There was a positional problem with the brush height today in the
outside brush station. When in the scrub position, the mechanical
stop wasn't making contact with the top of the micrometer. The
arm thats attached to the bimba cylinder seemed to have loosen
enough to cause this situation.
Therefore, the micrometers were checked to assure that the proper
settings still existed, which they did. So, we physically adjusted the
arm and re-tightened the the bolt. Checked for brush contact with the
wafer at the spin position. Tested fine.
A Shift
|
45.53 | Robot Fingers | FABSIX::R_GEE | | Mon Nov 13 1995 07:16 | 19 |
| * The robot fingers wouldn't close properly.
Checked the robot and spin station spindle for any loose
hardware, everything looked good.
Adjusted the teflon spring between the robot arms to even the
distance between the fingers and the wafer.
Changed the off-set position for Rotation at Spin from 871
to 879.
Changed the off-set position for Robot Fingers from 260
to 210.
Re-adjusted the Fingers Closed flag.
Ran 50 verification wafers, everything looks good.
B shift & Mitch
|
45.54 | DI attack!!!! | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Nov 23 1995 06:35 | 32 |
|
The status of SCRUB.A1 is as listed. DI water worked
its way through the bearings, through the spindle shaft
and the spindle tube. The bearings and tube are rusted and
are binding. The DI water also got down to the motor which
is why the machine went down in the first place. The motor
was stopping during the spin cycle. The motor is in stock
and should be replaced. Whom ever calls for the rush order,
see if they can assemble the outer spindle assembly. We
have limited access to any presses and would save alot of time.
I could put it back together but if OPS wants it when the parts
come in, it might be just as easy to get Ron in.
Parts needed to be ordered...
OnTrak ID #
(1) 13-0082-029 spindle shaft
(1) 13-0082-024 spindle tube
(1) 13-8082-073 splash gaurd, 200mm
(1) 90-0033-019 flat washer, 7/8 OD
(2) 201-036 ball bearing, .50 bore, 1.12
(1) 205-061 wave washer
In Stockroom 6:
===============
06-11110-01 - electrocraft motor
06-11078-01 - pulley (for motor)
BJ
|
45.55 | Motor replaced and old spindle installed | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Sat Nov 25 1995 05:19 | 12 |
|
I replaced the motor. I rebuilt the spindle and
shaft assembly. The bearings are shot and hummmm at
the high speed of 1800 RPM's which is the final step
of the spin process. The bearings will get hot and
won't last. The bearings also cause some extra vibration
in the tool. The entire spindle assembly is coming in
on either Monday or Tuesday. PLEASE use SCRUB.A1 as little
a possible. If both scrubbers are free, use #2.
BJ
|
45.56 | spin station blues | SUBPAC::LANDRY | | Sat Nov 25 1995 18:30 | 17 |
| The Spin Station shaft assembly was received in today. Due to
manufacturing constraints, the system wasn't relinquished to us.
Therefore, the assembly is located in the large compartment of the
grey roll-around tool box, should the system become available. By
the way, the assembly arrived fully assembled, not in the many pieces
we feared it would show up as.
Mid-afternoon, the system went down for a "robot fingers not closed"
error. One of the grippers on the robots left arm was making contact
with the spin station fingers. This was prohibiting it from picking
the wafer up properly. Adjusted the spin station lock position such
that the robot grippers clear the spin station fingers. After this
was achieved, the robot was "bouncing" at the wafer grip position.
Re-taught the grip position and the system ran fine.
Weekend warriors
|
45.57 | New Spindle installed, not quite up yet | SUBPAC::LANDRY | | Sun Nov 26 1995 19:17 | 27 |
| Today, the new spin station spindle assembly was installed in the system.
Many assembly learning curves along the way. Many post adjustments needed
by the end of installation. Probably not recommended for a one man job.
And still, the system is down, for no motor rotation.
After removing the old assembly, it was evident that the tools to properly
construct/install the assembly don't exist in CMP. Some hardships were
fitting the new bearing into it's retaining cavity (PVC pipe had to assist
on this one), Allen wrenches too tall, wrenches turning constraints, etc.
Anyway, everything finally came together mechanically. Had to adjust the
spindle height, had to adjust two sensors that apparently got knocked out
of position a bit, had to replace a broken micro-switch for the spin
fingers closed position. Those are all set now.
So, the current status is that the spin motor won't rotate. I see that the
inhibit switch is active, which I believe prevents the motor from receiving
power. Hopefully, there is info in the manual. I had talked with Ron
Simmons earlier in the evening. He says to look at a possible bad board
for the motor, but I hadn't noticed the inhibit LED on at the time. You
may want to consult him before you get too deep. He mentioned a PIA board
but reconfirm with him that this is the case.
Thanks go to Jesus Roa for his help with a lot of the installation. He was
extremely helpful.
Doc
|
45.58 | spin station spindle shaft | FABSIX::R_GEE | | Mon Nov 27 1995 07:04 | 17 |
| * Took over for A shift after the spin station shaft was replaced,
the motor wouldn't rotate.
- Checked the fuses and voltages, all o.k.
- Removed the drive belt from the motor and found the motor
was working fine.
- Noticed there was a mechanical bind in the spin station shaft.
Removed the shaft and found the bottom of the finger assembly
rubbing against the top of the bellows, adjusted and
re-assemblied.
- Calibrated the spin station spin speed and adjusted some
sensors, all set.
B shift
|
45.59 | rebuilt spin station shaft | STRATA::RIDLON | Eliminate the obvious | Tue Nov 28 1995 00:42 | 3 |
| Rebuilt the spin station shaft assembly with bearings found in stock.
Wanted to see if we could rebuild them without a press. Only took about
an hour and end result is as good as new.
|
45.1 | SCRUB.A1 INSTALL/MECHANICAL ACCEPTANCE | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:24 | 253 |
| SUBPAC::LANDRY 10 lines 6-JUN-1994 16:56
-< 6/6 update >-
--------------------------------------------------------------------------------
All,
KSI (the plumbing contractors) has been diligently working on the
installation of the exhaust/drain configuration. Yes, the drains and
the exhaust end up going out the same plumbing. What they have done
is run plumbing up to the system. They seemed to be moving along.
No other hookup activity was noticed today.
Doc
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.2 CMP-SCRUB.A1 2 of 145
SUBPAC::LANDRY 25 lines 16-JUN-1994 16:04
-< 6/16 Update >-
--------------------------------------------------------------------------------
Installation Update:
The GCW drain sections have been completed. They still need to tie
them into the exhaust. Said exhaust has yet to be run.
The Ammonium Hydroxide line has been installed.
The High Purity DI lines have been run.
The electrical is complete. They just had to supply a 120 VAC outlet.
The rack for the OCP's has been constructed and the OCP's have been
mounted. Awaiting hook up of the communication line to the main
system.
Earl Dawson and I brought up all the accessory equipment related to
this system. The "touch screen" monitor has been installed and one
of the smoked grey covers has been installed. Still need to attach
the last remaining cover. Will have to procure some hardware to
accomplish the feat.
The spare parts have been identified and stockroom personnel have full
control over them at this time.
Doc
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.3 CMP-SCRUB.A1 3 of 145
SUBPAC::LANDRY 16 lines 24-JUN-1994 14:12
-< 6/24 Installation Update >-
--------------------------------------------------------------------------------
* GCW drains complete
* NH4OH line complete
* UHPDI 75% complete
* Electrical complete
* Electrical for the OCP's is complete. Still need to have OCP comm
line to the system connected.
* Still waiting for exhaust work to begin
* Installed final smoked grey cover
Doc
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.4 CMP-SCRUB.A1 4 of 145
SUBPAC::TOOKER 24 lines 8-JUL-1994 11:03
-< Hook-Up Status For 7/8/94 >-
--------------------------------------------------------------------------------
CMP_SCRUB.A1 Tool Hook-Up Status For 7/8/94
COMPLETED ITEMS:
----------------
* DI Water connections
* Ammonium Hydroxide connections
* All GCW drain connections
* Eelectrical connections
OPEN ITEMS:
-----------
* CDA final connection
* All exhaust work
* OCP interface to system
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.5 CMP-SCRUB.A1 5 of 145
SUBPAC::TOOKER 14 lines 26-JUL-1994 09:20
-< ex >-
--------------------------------------------------------------------------------
CMP-SCRUB.A1 Tool Hook-Up Status As Of 7/25/94
PUNCH LIST ITEMS
----------------
ITEM RESPONSIBLE PERSON
--------------------------------------------------------------------------
1) Complete exhaust work. D. Atkinson/NEEVCO
2) Install additional pipe d. Atkinson/NEEVCO
support on exhaust.
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.6 CMP-SCRUB.A1 6 of 145
SUBPAC::LANDRY 9 lines 1-AUG-1994 13:14
-< 8/1 equipment installation update >-
--------------------------------------------------------------------------------
Brian and I ran the tubing and hooked up the CDA to the system. This
completes all the plumbing hookups to the system. The only remaining
hook up concerns the communication connection between the main system
and the OCP.
We also repaired the magnets on the plexiglass access doors. The
magnets had become unglued from the plexi........so we reglued them!
Doc
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.7 CMP-SCRUB.A1 7 of 145
SUBPAC::BJUBINVILLE 16 lines 2-AUG-1994 16:42
-< Ron Simmons has arrived!! >-
--------------------------------------------------------------------------------
Ron Simmons was in this afternoon so here's the scoop.
The facility guys were turning on the N2, DI water and CDA when
they discovered some leaks near CMP.A1 so the scrubbers were
on hold. Ron will be back in the morning around 8am to start
things into purge. The power is all hooked up but the 15A
breaker is off on both machines. There was a concern with a
fitting on the 1/2" ammonium line. Ron Simmons would rather
not use stainless. He doesn't need the fitting to put the
machine into purge so he'll bring the fitting/parts with him
on Monday.
BJ
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.9 CMP-SCRUB.A1 9 of 145
SUBPAC::BJUBINVILLE 11 lines 9-AUG-1994 16:02
-< we're getting there >-
--------------------------------------------------------------------------------
Ron I and worked with the wafers on this scrubber
today. There was a screw on the spin station carriage
plate that must of broken off during shipping. We also
realigned the robot transfer arm for a smoother pick-up.
The brushes were installed as well. We are going to purge
the NH4OH with DI in the morning. We cycled through (50)
wafers without any problems.
BJ
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.10 CMP-SCRUB.A1 10 of 145
SUBPAC::BJUBINVILLE 10 lines 10-AUG-1994 16:14
-< 8/10 passdown >-
--------------------------------------------------------------------------------
The guys downstairs wanted the ammonia line flushed with DI
water for 24 hours and the pressurized N2 after that. I have
this scrubbed in full purge as of right now. I'm take a 1/2
day tomorrow so I'll be filling Doc in on where and when with
the N2. Ron came in this morning and tied up some loose ends
and to make everything was ok, which it was.
BJ
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.11 CMP-SCRUB.A1 11 of 145
SUBPAC::LANDRY 12 lines 11-AUG-1994 19:08
-< purged and purified!!! >-
--------------------------------------------------------------------------------
The H20 Purge on the ammonium Hydroxide line was completed at 3:00 PM.
The N2 purge on the line was then started and completed. The NH4OH
line was then re-attached to it's proper port. We could possibly see
NH4OH to the tool as soon as Friday.
We still need to have the OCP's interfaced to the main system
We attempted to run the systems today with a "burn-in" recipe. The
system expirienced several timeout/sensor problems. The rep will
be in to address the systems on Friday.
Doc
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.18 CMP-SCRUB.A1 18 of 145
SUBPAC::BJUBINVILLE 23 lines 22-AUG-1994 17:01
-< Passdown for 8/22 >-
--------------------------------------------------------------------------------
The problem with the outter brush seems to have been fixed
after Ron replaced the amplifier and A2 PIA chip. The machine
is back on track with 150 wafers through it towards its 1000 wafer
goal.
The OCP interface that was designed for us through OnTrak
seems to be wrong. Their schematic seems to have been misplaced
so another one is being faxed to us tomorrow. We will be evaluating
this further tomorrow. Jim T. will be looking the original specs
that DEC layed out to OnTrak for our request.
The ammonia needed to be regulated at 35psi. but it was never
done. Dan Smith is looking to get me a teflon regulator and teflon
gauge in the morning. Unregulated, it will be around 45-50psi.
The DI water pressure was still sitting around 30psi at 3:30pm
today. Bob Kerwin did come into CMP to look at the gauge but it
remains to be unseen if anything was accomplished beyond 3:30pm.
BJ
|
45.2 | COMPUTER INITIALIZATION PROBLEMS | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:24 | 38 |
| SUBPAC::BJUBINVILLE 35 lines 15-AUG-1994 18:07
-< Passdown for 8/15 >-
--------------------------------------------------------------------------------
* Ron and I came into where the SCRUB_A2 was stopped and had
homing problems. After we rebooted the machine, there was
a problem on boot up. The machine is suppose to go right
into the OnTrak self initializing mode but it was staying
at the WinShell prompt. After Ron called the software
guru from OnTrack he had use copy a file over from another
directory. These machines are touch screens so we had to hunt
down a keyboard with a RS232 connector. We had to type the
following as these machines are not DOS operated.
WinShell:> COPY DATA2 DATA -R
After this is typed, reboot the machine it ok. Every 2 minutes
the machine does a self back-up of essential data files that usually
takes a couple of milliseconds. If you get acurate enough, if you
shut the machine the exact same time as this back-up starts, the
machine will put all of the robot offsets reset to factory defaults.
OnTrack said this has only happened twice in our model history.
This is very rare. I copied the offsets for both machines into
their respective logbooks. I will eventually copy them into here.
* The DI water is still an issue. I spoke with Bob Kerwin and
the issude is still being addressed. He's looking at least
another week before we some action.
* Repaired a few more leaks
* we went over some maintainence training as well as some process
issues with Steve.
BJ
|
45.3 | FLOW VALVE WON'T ADJUST HIGHER, MOTOR SPEED OUT OF SPEC | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:26 | 22 |
| SUBPAC::BJUBINVILLE 23 lines 16-AUG-1994 13:10
-< Passdown for 8/16 >-
--------------------------------------------------------------------------------
We calibrated the motor speed offsets to match the
"Burn-In" recipe.
We found the outter bottom brush flow valve not reaching
over 400ml, it should of been higher. We replaced the valve and
found a defect in the valve. We borrowed on from the Wet-Lab
and got this this one going. Ron will bring the borrowed parts
to get the one in the Wet-Lab ready.
The machine has continued to be in mechanical acceptance
and has gotten up to around 700 wafers now.
BJ
This machine is still in the 1000 maintainence
acceptance.
BJ
|
45.4 | SENSORS AND MOTOR SPEED PROBLEMS | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:27 | 50 |
| SUBPAC::BJUBINVILLE 24 lines 18-AUG-1994 16:24
-< Passdown for 8/18 >-
--------------------------------------------------------------------------------
* We finally ironed out the bugs. The was sensor problem
with the inside brush station. The alarm kept tripping
because the sensor would detect a wafer but would stay
on instead of shutting off. We traced it back to a 3v
drop across the sensor/amplifier. We replaced the IC
chip for this sensor and the voltage got back up to 5v.
* We set the brush height to 14.5mm pending results from
Tencor. We might have to either raise or lower them.
* We tweaked in the outside, inside and spin motor speed
offsets to stablize them.
* The receiver door switch had the same problem as #2 yesterday.
The LOCKED and UNLOCKED lights stayed lit when only the LOCKED
was suppose to. We had to adjust the sensor.
We've run about 150 wafers straight with now problems. Keep
your fingers crossed.
BJ
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.15 CMP-SCRUB.A1 15 of 145
SUBPAC::BJUBINVILLE 15 lines 17-AUG-1994 17:27
-< passdown for 8/17 >-
--------------------------------------------------------------------------------
The initial problem was that the robot wasn't picking
up the wafer properly. We retaught the robot position to
get in.
A "inside brush motor speed out of spec" error was coming up
all the time. We traced it to the drive belts on the rollers.
We tightened them up and the motor feed back was stable but just
a few RPM's under 40. We are going to tweak it to 40 tomorrow.
There also seems to be a problem with the inside brush sensor. Ron
replaced the PIA #1 board to try and get it to behaive.
BJ
BJ
|
45.5 | SEALS REPLACED, LEAKING MANIFOLDS AND MOTOR SPEEDS | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:28 | 18 |
| SUBPAC::BJUBINVILLE 15 lines 25-AUG-1994 17:02
-< Passdown for 8/24 >-
--------------------------------------------------------------------------------
* Inner brush rear roller shaft was leaking, replaced
seal
* Outer brush rear roller shaft was leaking, replaced
seal
* Outer exit manifold leaking out end (brush lift end),
replaced manifold.
* tweaked in inside, outside and spin motor speed offsets
BJ
|
45.6 | CARRIAGE PROBLEMS | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:28 | 26 |
| SUBPAC::LANDRY 22 lines 20-SEP-1994 14:53
-< Get a "grip" on it man! >-
--------------------------------------------------------------------------------
Today..........
* The system shattered a wafer. First off was to clean the big mess
it made. Picked up as many pieces as possible by hand and disposed
of them in the proper recepticle. Removed the remaining fragments
by use of the hose vacuum. Wiped down affected surfaces.
* Traced the problem to be with the retracting roller assembly NOT
being level. As the grippers came in to hold the wafer, it didn't
secure it properly. Of course I didn't see it actually happen or I
would have stopped it before it broke the wafer. But, through
monitoring it's performance on the next several wafers, I was able
to see this almost happen again. Found the roller assembly bracket
to be loose.
* Jim T. and myself then manually operated the system to align the
roller assembly properly and then secured the bracket back to it's
optimum operating setting.
Doc
|
45.7 | AMMONIA TURNED ON | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:29 | 21 |
| SUBPAC::LANDRY 18 lines 30-SEP-1994 14:29
-< Is that a babies diaper? No, thats the NH4OH! >-
--------------------------------------------------------------------------------
* AMMONIA IS FLOWING!
* The lockout/tagout has been removed from the 2% NH4OH line.
* We (Jim Sadin and myself) edited a program to get the NH4OH to flow.
The internal shut off valve (or turn on valve to some) was opened,
the flow meters for were opened/adjusted, and the lines were
observed for leaks (no leaks found).
* We did a quick D.O.E. for NH4OH flow with both systems A1 & A2
running at the same time. We were capable of attaining flows of
7-8 on the flow meter. Both systems fluctuated periodically.
* The programs were returned to their normal settings for NH4OH,
which is 20 seconds for the outside brush and 10 seconds for the
inside brush.
Doc
|
45.8 | SOFTWARE UPGRADE | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:30 | 16 |
| SUBPAC::BJUBINVILLE "Quit crying and play hockey!" 13 lines 7-OCT-1994 15:58
-< Passdown for 10/7 >-
--------------------------------------------------------------------------------
Ron came in and completed yet another software
upgrade to this machine. It was designed to
eliminate some small bugs that were irrevalent
to us but none the less, they're gone. The
new rev. # is listed in the book.
Ron also replaced the (4) spin station fingers
to a bigger style with better wafer handling.
Wafers were run through without any problems.
BJ
|
45.9 | DOUBLE STACKED WAFER | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:31 | 20 |
| <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
SUBPAC::BJUBINVILLE "Quit crying and play hockey!" 14 lines 13-OCT-1994 16:09
-< Passdown for 10/13 >-
--------------------------------------------------------------------------------
The RECEIVER access door was opened during the run causing
the machine to stop. There was a wafer in the otter brush
that wasn't recovered when someone RESET the station. The
next wafer through piggy backed through the rest of the
cycle. If you want to reset a station and there is a wafer
in that station, recover the wafer FIRST. If you don't,
when you RESET the station, you are just telling the machine
there is now no wafer in the station when there really is.
Thus.... piggy back.....
BJ
|
45.10 | HEAT LAMP NOT COMING ON | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:31 | 12 |
| SUBPAC::BJUBINVILLE "Quit crying and play hockey!" 9 lines 2-NOV-1994 16:23
-< Passdown for 11/2 >-
--------------------------------------------------------------------------------
The lamp wasn't coming during its cycle. The lamp
was in the socket but wasn't making good contact. I
believe it was due to vibration in the room. I will
start to look into getting some kind of a clip to hold
it into place.
BJ
|
45.11 | BRUSH CHANGE | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:31 | 23 |
| SUBPAC::LANDRY 20 lines 8-NOV-1994 16:47
-< Bald people shouldn't have to do brushes!!!! >-
--------------------------------------------------------------------------------
Todays activities.......
What a day! Ron Simmons came in today to assist in changing out the
PVA brushes on the system. We started about 10:30 AM and just finished
up at 4:30 PM. We definitely had some big learning curves! Let me
explain.
First off, the brushes were extremely difficult to install onto their
cores. The PVA material is easily stretched and the studs are easy
damage. The brush must be put onto the core evenly. This proved to be
the biggest problem with the installation. Bobby, Steve and myself all
to a crack at putting these brushes onto the core. I'd say all three
of us didn't fair to well. Although the brushes were installed to what
we thought to be pretty good, we were proven wrong upon core
installation. The peaks and valleys were not evident until the brushes
/cores were rotated. So off they came again to be retried. Finally,
many hours later, the brushes were installed properly. Steve will now
begin to perform particle checks. Lets hope all is well.
Doc / Bobby
|
45.12 | MEGASONIC SWEEP | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:32 | 13 |
| SUBPAC::BJUBINVILLE "Quit crying and play hockey!" 10 lines 10-NOV-1994 11:31
-< Passdown for 11/10 >-
--------------------------------------------------------------------------------
Engineering was inquiring on the dwell time
on the megasonic sweep. We looked into the LOW
and HIGH flow sensor settings and the window was
smaller than A2. We brought them to the same setting.
Another thing was the megasonic flow was at 800ml.
We matched it at 900ml like A2 as well. The dwell is
very minimal and acceptable at this time by Engineering.
1ft shift boys
|
45.13 | WAFER ROLLER NOT SPINNING | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:32 | 10 |
| ASDG::POIRIER 7 lines 22-NOV-1994 23:25
-< Guide rollers in brush station #2 >-
--------------------------------------------------------------------------------
* Jim Sadin noticed that the wafers were not spinning in brush station #2.
Bob and I tightened the set screws on the guide rollers and the
wafers now spin normally. It appears that the rollers were slipping on
the drive shafts.
John.
|
45.14 | | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:32 | 8 |
| PATE::GEE 5 lines 28-NOV-1994 23:35
-< Outside station lower brush >-
--------------------------------------------------------------------------------
The outside station lower brush had some hesitation as it rotated in
idle speed. Installed a missing bearing that sits inside the brush
core end from stock.
Bob
|
45.15 | ROBOT PROBLEMS | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:33 | 17 |
| SUBPAC::LANDRY 14 lines 1-DEC-1994 17:02
-< Who needs a wafer dicing saw? >-
--------------------------------------------------------------------------------
Today, the robot malfunctioned, breaking a wafer (exercise wafer) very
neatly in two half moon pieces. It broke so cleanly that one half
actually stayed positioned in the robot fingers. How it broke is a bit
sketchy! The robot picked up the wafer from the spin gripper. It is
believed that the robot didn't properly rotate 90 degrees prior to
attempting it's move to the unload station. The robot was found with
it's finger (be nice here) up against the plexiglass cover.
The system was EMO'd when the breakage occured (good quick judgement).
Upon repowering the system, all stations were initialized. Ran several
tests wafers to confirm problem had vanished. We'll have to keep an
eye on this for a bit.
Doc
|
45.16 | INSIDE BRUSH TIMEOUT ERROR | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:34 | 15 |
| SUBPAC::BJUBINVILLE "Quit crying and play hockey!" 12 lines 7-DEC-1994 14:32
-< Passdown for 12/7 >-
--------------------------------------------------------------------------------
The scrubber kept going into CLEARING after every
three wafers. Machine would alarm and say the scrubber
timed out looking for the wafer at the inside brush
station. If an error involving any station timeouts
or lost/broken wafer, there's a good chance you'll
need to RESET the receiver. When you reset the receiver,
the computer re-sync's the counters which gets all of the
stations on the same time plane.
BJ
|
45.17 | ROBOT PROBLEMS | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:34 | 16 |
| SUBPAC::BJUBINVILLE "Quit crying and play hockey!" 13 lines 9-DEC-1994 11:46
-< Passdown for 12/9 >-
--------------------------------------------------------------------------------
Ron came in for a visit. I told him about the
two wafers that broken in the last two weeks. We
went over the robot and spin station and found some
problems. When the spin station first grabs the wafer,
it closes twice to take the wafer. We found there was
lack of air pressure to take the wafer on all four
fingers. After we adjusted the air to the fingers,
it grabbed fine. The robot fingers weren't closing
all the way either, same fix. We bent the tabs in
on the lamp socket so the lamp will hold tighter.
BJ
|
45.18 | SQEALING NOISE, BRUSH ROTATION HESITATION | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:35 | 26 |
| <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.54 CMP-SCRUB.A1 54 of 145
SUBPAC::BJUBINVILLE "Quit crying and play hockey!" 19 lines 3-JAN-1995 07:02
-< Passdown for 1/3 >-
--------------------------------------------------------------------------------
I went through with some things with Bob concerning the
squealing noise coming from the upper brush supporting mounts.
We re-aligned the supports to get the noise to go away. Bob
also noticed that the lower brush on both stations seemed to
be stopping and hesitating. We traced it to the front bearings
on both brushes. We replaced them and everything ran fine. There
are no more bearings in stock. I put in a shortage request as well
as up-ing the min/max level to 16 bearings.
BJ
|
45.19 | BLADE LIFTER NOT REPEATING | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:37 | 39 |
| SUBPAC::BJUBINVILLE "IT'S HOCKEY TIME!!" 10 lines 2-FEB-1995 09:52
-< Split Ends, isn't that a hair thing? >-
--------------------------------------------------------------------------------
This morning, Ron Simmons came to look at our
blade lifter issue. It was traced to the drive chain
for the lifter. In one section, it was split up the middle
for about an inch. The split in the belt during the travel
wasn't as effective as when the split got the bottom pulley,
it was right when the blade was at its top position. This
is where it was not repeating. The belt is getting replaced
within the next half hour.
BJ and Doc
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.64 CMP-SCRUB.A1 64 of 145
SUBPAC::LANDRY 18 lines 1-FEB-1995 18:31
-< Whats up with the Knife? >-
--------------------------------------------------------------------------------
The system is currently down for a robot to knife transfer problem.
The blade isn't returning to it's proper receive position everytime.
The cassette indexing isn't consistant either. It has double loaded
wafers in the same cassette slot. This means that the cassette index-
ing is moving enough to clear the inhibit position, but not far enough
to make it to the next slot. Once this happens, the cassette homes,
even though there are more wafers coming through the line. Bizarre!
The problem appears to be a PC board problem. Bobby Gee and I swapped
out the Intellegent Stepper/Driver board with one from stock. This
board controls both the knife and cassette indexing. No luck. Next
most possible culprit is the PIA board. This board, once installed,
will require calibration. Possible mechanical problems with the unload
station are remote, but still possible. Good Luck.
Ron Simmons can be in tomorrow morning if needed. He will call us in
the AM.
Doc
|
45.20 | BRUSHES WON'T ROTATE IN IDLE | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:37 | 9 |
| PATE::GEE 6 lines 2-FEB-1995 23:21
-< Scrub Station >-
--------------------------------------------------------------------------------
The brushes would not rotate in Idle mode. Found that the
wrong process program was loaded. Loaded the correct program
"STRAS_CLEAN_BEST", System is working fine.
John & Bob
|
45.21 | | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:37 | 204 |
| SUBPAC::BJUBINVILLE "IT'S HOCKEY TIME!!" 201 lines 6-FEB-1995 01:48
-< Machine Characterization as of 2/6/95 >-
--------------------------------------------------------------------------------
CMP_SCRUB.A1
ONTRAK DSS-200 SCRUBBER
MACHINE MAINTENANCE CHARACTERIZATION SHEET
LOAD AREA
==========
____
ELEVATOR: PHYSICAL POSITION ||_| <-- (1)
(measurement taken from (1)top || ^
of the elevator mounting || ^
bracket to the (2)bottom of the || | <-- (2)
motor mount casing. ||||_____ <--elevator base
UPPER POSITION (SLOT #1) -- 1 27/32"
HOME POSITION -- 1 3/16"
LOWER POSITION (SLOT #25) -- 7 13/16"
MACHINE OFFSET -- (found in ROBOT OFFSETS menu)
LOAD STATION -- 1310
SPRAY MAINIFOLD --
(bar and nozzle position) verify alignment of scribed marks
============================================================================
SCRUB MODULES
=============
MICROMETER SETTINGS: verify ENGINEERING set point
INSIDE BRUSH STATION | OUTSIDE BRUSH STATION
SETTING/ACTUAL | SETTING/ACTUAL
--------------------------------------------
|
WAFER ROTATION (TACH READING IN RPM'S) 40 / 44 | 50 / 49.5
|
BRUSH ROTATION (TACH READING IN RPM'S) 118 / 118 | 118 / 121
SPRAY MAINIFOLD --
(bar and nozzle position) verify alignment of scribed marks
==============================================================================
SPIN STAION
===========
SETTING/ACTUAL
--------------
WAFER ROTATION (TACH READING IN RPM'S) 1800 / 1715
500 / 569
CUP HEIGHT -- (measurement taken UPPER POSITION -- 7 9/16"
from very top of any
finger to bottom of MIDDLE POSITION -- 6 15/16"
tub)
LOWER POSITION -- 4 7/16"
GRIPPER FINGERS -- (measurement taken FINGERS OPEN -- 8 1/2"
from center of left
rear finger to center FINGERS CLOSED -- 7 1/2"
of left front finger)
WAFER TEMPERATURE -- AS WAFER ENTERS SPIN STATION -- 75 F
AS LAMP IS ON -- 80 F
AFTER LAMP SHUTS OFF -- 75 F
EXHAUST VELOCITY (in inches) -- 1.3
============================================================================
ROBOT
=====
OFFSETS -- WRIST AT SPIN STATION -- 108
WRIST AT RECEIVE -- 240
SPIN XPORT DELAY -- 40
ROTATION AT SPIN STATION -- 882
ROTATION AT RECEIVE -- 1813
ROBOT FINGERS -- 370
ROBOT FINGERS -- (measurement taken OPEN -- 9 7/16"
from O.D. to O.D.
from finger to finger CLOSED -- 8 1/4"
==========================================================================
UNLOAD MODULE
=============
CASSETTE AT POSTION SLOT #1-- (insert new cassette
and let machine go to
position and slot #1. -- 16 3/4"
measure from the outer
most part of the back
panel the middle of slot
#1)
KNIFE EDGE -- (measure from very UPPER POSITION -- 10 1/4"
top of knife to the
top of panel that the LOWER POSITION -- 7/8"
edge goes into)
MACHINE OFFSET -- (found in ROBOT OFFSETS menu)
UNLOAD STATION -- 490
BLADE LIFTER -- 11500
=============================================================================
ELECTRONICS
===========
IDLE | RUNNING
========================
POWER SUPPLY VOLTAGE --
+5 VOLT POWER SUPPLY 5.15 | 5.15
|
+12 VOLT POWER SUPPLY 12.3 | 12.37
|
-12 VOLT POWER SUPPLY -12.38 | -12.38
|
+24 VOLT POWER SUPPLY 24 | 24
|
+48 VOLT POWER SUPPLY 48.2 | 48.2
SERVO DRIVE POWER SUPPLY-
+40 VOLT PICK-OFF 41 | 41
+5 VOLT PICK-OFF 5 | 5
============================================================================
FLOW SETTINGS
=============
SEND -- 5
AMMONIA OUTSIDE STATION -- 3
OUTSIDE BRUSH UPPER -- 10
OUTSIDE BRUSH LOWER -- 7
AMMONIA INSIDE STATION -- 3
INSIDE BRUSH UPPER -- 9
INSIDE BRUSH LOWER -- 7
SPIN STATION UPPER -- 3
SPIN STATION LOWER -- 3
MEGASONIC NOZZLE -- 9
===========================================================================
PRESSURE SETTING 50 PSI
================
SIGNET FLOW SETTING
=================== LOW -- 40
HIGH -- 0
|
45.22 | | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:37 | 15 |
| SUBPAC::LANDRY 12 lines 13-FEB-1995 17:11
-< New Software installed. >-
--------------------------------------------------------------------------------
Ron Simmons came in a started performing a couple of up-grades to the
system. New software was installed and checked out. The new software
will protect all data entry regardless of when the system is shutdown.
The old version didn't have this feature and it took several hours to
re-enter all the datat once it was lost, ask Brian, he'll tell you what
he went through the time this happened to us.
Ron has a couple of more items to perform when he and the system are
both available.
Doc
|
45.23 | coiled electrical cord for the send station elevator | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:38 | 39 |
| SUBPAC::LANDRY 12 lines 22-FEB-1995 16:35
-< Minor problem needs to be addressed w/ Ontrak >-
--------------------------------------------------------------------------------
The coiled electrical cord for the send station elevator has been
chaffed conciderably. It was beginning to create particles on the
outside of the plexiglass box (no contamination within the box where
the wafers are loaded). As a temporary provision to stop this from
happening, a strand of electrical tape was applied.
OnTrak will be notified of the problem. We'll get the cords replaced
as well as possible changing the "style" of hardware that protrudes
from the plexiglass box. It currently has hex-head hardware where a
button head style would probably work out better. Stay tuned!
Doc
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.86 CMP-SCRUB.A1 86 of 145
SUBPAC::LANDRY 16 lines 7-APR-1995 14:14
-< wire harness replaced >-
--------------------------------------------------------------------------------
Ron Simmons, OnTrak Rep Extraodinaire, came in today to replace the
wiring for the Load Staion elevator power. The original wiring was
mounted such that the wiring would rub against the hardware protruding
from the load station tub. The new wiring has been installed such that
it will no longer rub in the previous manner. The socket head screws
will be replaced in the near future to button head screws, hence
preventing this chaffing to occur again.
We also checked out the possible leak on the feedthroughs in the brush
stations. Water was noticed to have formed there on one of his
previous visits. The last two times he was here, they have been dry,
as should be. Original sighting could have possible been attributed to
wet transfer boats, as they are some times temporarily placed there.
We'll monitor for future sightings.
Doc
|
45.24 | HOT DI WATER TIE IN | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:38 | 16 |
| SUBPAC::LANDRY 13 lines 23-MAR-1995 17:07
-< Sometimes its hot sometimes its not! >-
--------------------------------------------------------------------------------
Heres what we got.......
The Hot DI water loop hookup has been performed. The system
now has the capability to dispense hot or cold water. Currently,
the cold (not really cold, but the regular di water normally going
there) is turned on to the system. The Hot loop's valve has been
locked out / tagged until it is ready to use. Each input line is
equipped with a temperature gauge.
The status of the equipment at this time remains as VERIFY, on
account that we gotta make sure the particle counts are acceptable.
1st Shifters
|
45.25 | | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:39 | 17 |
| SUBPAC::LANDRY 14 lines 24-MAR-1995 17:29
-< Hot DI Water Loop Status >-
--------------------------------------------------------------------------------
The lastest scoop on the Hot Di Hookup goes like this.........
We coordinated some work today with Holden Mechanical. They had
introduces water into the Hot DI line, when in fact, we weren't ready
to receive it. Therefore, we had them disconnect a fitting in the Sub
Fab to drain the line to the system. Then Roger and I hooked up an
N2 purge to the line to dry it out. After many hours, Holden checked
the line at their end. They still found traces of moisture. Therefore,
it was decided that we would leave the N2 purge on until the line was
completely dry. Holden will periodically check the status and let us
know when all is dry.
Doc / Roger
|
45.26 | | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:39 | 9 |
| ASDG::COOPERMAN 6 lines 25-APR-1995 13:22
-< water flow adjustment >-
--------------------------------------------------------------------------------
Adjusted flow for top brush on .a1 to 9 from ~5-6. Should
have been at 9, not sure why was low.
Steve
|
45.27 | | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:39 | 12 |
| SUBPAC::BJUBINVILLE "NFD Fire - Rescue" 9 lines 26-APR-1995 10:50
-< PVA brushes replaced and being aligned >-
--------------------------------------------------------------------------------
Machine ended up failing MQC's for high particles. It
was either leave it down or try changing them. After they were
changed, the scrubber came right into spec. However.... they
weren't perfectly asymmetrical. With Larry and John, we gave
it another shot. They are coming in nice. They should be finishing
up real soon.
BJ
|
45.28 | | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:39 | 9 |
| SUBPAC::BJUBINVILLE "NFD Fire - Rescue" 6 lines 12-MAY-1995 06:23
-< Roller speed out of spec >-
--------------------------------------------------------------------------------
Came to a OUTSIDE brush roller out of spec. This error
indicates the rollers, not the brushes. A recalibration was
performed and the rollers are running fine now without any
machine errors.
BJ
|
45.29 | | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:39 | 14 |
| FABSIX::R_POIRIER 11 lines 18-MAY-1995 23:52
-< Megasonics Sweep Problem >-
--------------------------------------------------------------------------------
There was a problem with the Megasonics Sweep home sensor. We
replaced the Optical Sensor (DEC P/N 94-072339-01) from S/R 94
and Megasonics Sweep Home works fine. Ran 26 dummy wafers to
check out spin station. Particle checks should be run before
using CMP-SCRUB.A1 for product.
2nd Shift
|
45.30 | | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:39 | 13 |
| SUBPAC::BJUBINVILLE "NFD Fire - Rescue" 10 lines 30-MAY-1995 07:56
-< ROBOT TIMING OUT >-
--------------------------------------------------------------------------------
The system kept alarming because the sensor for FINGERS
CLOSED wasn't being seen. After powering down the system
and removing the covers to the robot, everything was in
order and working fine. A dozen or so wafers were run on
the machine without any problems. About an hour later, the
problem came back but was placed in WAFER RECOVERY before I
was able to see the problem. The machine ran fine since.
BJ
|
45.31 | | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 01 1995 23:39 | 13 |
|
SUBPAC::BJUBINVILLE "NFD Fire - Rescue" 8 lines 23-MAY-1995 06:18
-< Megasonic arm linkage >-
--------------------------------------------------------------------------------
The megasonic arm linkage was all out of whack. After the
arm was repositioned and the amplifier adjusted it moved fine.
A valve and tee were installed on the DI line for the DI gun
that will be used only for brush changes.
BJ
|
45.60 | ROBOT OFFSETS CHANGED | FABSIX::B_WESTPHAL | | Sat Dec 02 1995 17:06 | 7 |
| SCRUB.A1 - Down for robot not picking wafers from the spin station. Found that
the robot fingers were closing improperly around the wafer.
Adjusted the level and positioning of the spin station by adjusting
the 4 base bolts. Also made several adjustments to the WRIST AT
SPIN, ROTATION AT SPIN, and FINGERS (120,868,210) offsets, ending up
with the final values of 124, 873, and 190 respectively. Ran a full
boat of wafers w/o any problems and returned the tool to production.
|
45.32 | ROBOT TIMING OUT | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:00 | 13 |
| SUBPAC::BJUBINVILLE "NFD Fire - Rescue" 10 lines 30-MAY-1995 07:56
-< ROBOT TIMING OUT >-
--------------------------------------------------------------------------------
The system kept alarming because the sensor for FINGERS
CLOSED wasn't being seen. After powering down the system
and removing the covers to the robot, everything was in
order and working fine. A dozen or so wafers were run on
the machine without any problems. About an hour later, the
problem came back but was placed in WAFER RECOVERY before I
was able to see the problem. The machine ran fine since.
BJ
|
45.33 | RECEIVE TIMEOUTS, BRUSH STATION TIMEOUT | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:01 | 20 |
| SUBPAC::LANDRY 17 lines 30-MAY-1995 18:27
-< National Robot Month continues on >-
--------------------------------------------------------------------------------
A couple of occurances today that required some attention...........
* With a hot lot running, I was summoned in to look at the system.
Seems the robot picked up the wafer, then just stalled out. The
alarm screen said that the was a timeout looking for the receiver
cassette. There was a receiver in there and ready to receive a
wafer, so I hit start and the run finished fine. Seems like someone
may have interrupted the run to fetch some completed wafers in the
receiver???
* Then there was a brush station timeout. Again, nothing was found to
be wrong. Diagnostically reset the station, but couldn't get the
load cassette transport belt to turn on. Had to reboot the software
to return to normal.
Doc
|
45.34 | SYSTEM DOUBLE LOADING | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:01 | 11 |
| SUBPAC::BJUBINVILLE "NFD Fire - Rescue" 8 lines 31-MAY-1995 04:29
-< double loading a spin station >-
--------------------------------------------------------------------------------
Two wafer got into the spin station at the same time
for some unknow reason. Machine had to bew powered down
for safe removal of the product wafers. A boat of wafers
was run without and problems.
BJ
|
45.35 | WAFER ROLLER ARMS NOT OPENING | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:02 | 15 |
| SUBPAC::LANDRY 12 lines 31-MAY-1995 17:55
-< Open up them arms! >-
--------------------------------------------------------------------------------
The system had some timeouts again today...........
Timeout looking for inside brush wafer. Problem actually is with the
outside brush station. When it's time to release the wafer from the
outside station, the wafer rotation arms were not retracting everytime.
The bimba shafts for this movement were cleaned and very lightly lube'd
with a drop of oil (no worry, the bimbas are located outside of the
brush stations and pose no contamination threat). The problem has not
occured since the lube was placed.
Doc
|
45.36 | RECEIVE CASSETTE NOT READY, SENSOR ADJUSTMENT | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:03 | 18 |
| SUBPAC::LANDRY 15 lines 1-JUN-1995 17:54
-< sensor adjustment had to be made >-
--------------------------------------------------------------------------------
The system came up with a "receiver not ready" (or something close to
that terminology) error today. Well, there indeed was a cassette in-
place and it was ready to receive a wafer. Went into the I/O functions
and found that the cassette present sensor was not being made. Looked
at the sensor arraingement and found out that the sensor is adjusted by
a set screw mounted into the arm above the sensor. The set screw was
very loose in the threads of the hole. This caused the screw to move
up the threads and out of position. Adjusted said switch and tested
results. All is fine for now.
With the looseness of the set screw, I would not be surprised to see
this sensor adjustment made again in the future. Stay tuned!
Doc
|
45.37 | ROBOT ARM PROBLEM | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:04 | 13 |
| FABSIX::R_POIRIER 10 lines 15-JUN-1995 23:43
-< Robot arm problem >-
--------------------------------------------------------------------------------
Robot arm kept stopping after picking up a wafer from the
spin station but would not put it in the cassette. This
problem was not constant. Brian found that the load cassette
switch was not being activated all the time when the cassette
was put in. The set screw had loosened and wasn't making contact.
Tightened the set screw and the problem went away.
|
45.38 | ROBOT FINGER PROBLEMS | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:06 | 43 |
| <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.118 CMP-SCRUB.A1 118 of 145
SUBPAC::LANDRY 36 lines 29-JUN-1995 19:46
-< Robot Positioning Problem >-
--------------------------------------------------------------------------------
An educational day..............
The system robot functions were messed up this morning. The
fingers were not closing to their proper position, causing the system
to "timeout" for finger position.
Found the several problems with the system. Problem one was the
flags for the open + closed potitions were out of whack. For this fix,
the open flag must be set up first. Once done, the motor steps take
over from there. The motor steps dictate the position of the closed
position. The mechanical flags are only used as indicator for the
diagnostic menu's. The closed flag mechanism was then moved to it's
proper location.
After this was setup, we attempted to run some wafers. We found
the gripper arms to be over traveling, causing the grippers to flex
back. The closed position was re-taught. Transfered several wafers
fine, but occasionally, it would mis-grip the wafer. Upon checkout,
we found that the "rotation at spin station" and the "wrist at spin"
positions were out of tolerence. These were causing the gripper
fingers to pick the wafer in the wrong spot on the gripper fingers.
By this I mean that the wafer was not aligned with the slots on the
grippers, but instead it was aligned with the shaft. After the
positions were re-setup, the system was transfering quite nicely.
T%he current System Offset settings are:
Load Station = 1310 Rotation @ Spin Station = 877
Unload Station= 490 Rotation @ Receive Sta = 1813
Wrist @ Spin Sta = 114 Balde = 11500
Wrist @ Receive = 240 Robot Fingers = 420
Spin Xport Delay = 40
Doc / Jim T
|
45.39 | BRUSH ROLLER ARM CRACKED | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:07 | 10 |
| SUBPAC::BJUBINVILLE "NFD Fire - Rescue" 7 lines 19-JUL-1995 06:33
-< Roller arm cracked >-
--------------------------------------------------------------------------------
Wafers kept sliding by the outside brush station and
doubling up in the inside brush station. The right roller
arm was cracked and causing the roller not to close to stop
the wafers. The arm was replaced and aligned.
Bob, Leon, BJ
|
45.40 | WAFER ROLLER ARMS NOT OPENING | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:08 | 20 |
| <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.122 CMP-SCRUB.A1 122 of 145
SUBPAC::LANDRY 13 lines 19-JUL-1995 20:10
-< roll them arms back >-
--------------------------------------------------------------------------------
Twice today, the system faulted for a timeout, waiting for a wafer to
release from the outside brush station. We found the problem to be
intermittent. After several test wafers, the roller arms didn't open
to allow the wafer to move on to the next station. The Bimba cylander
appeared to be a bit sticky. Cleaned the shaft and lightly lubricated
using light oil. Ran several test wafers and it appeared to be smooth.
An hour later, the problem happened again. Began to run test wafers.
No problems occurred. After 40+ wafers were run through it, the status
was changed to UP. We'll have to monitor it to see if further problems
arise.
Doc
|
45.41 | COMPUTER FREEZE UP, ROBOT FINGER PROBLEM | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:09 | 24 |
| SUBPAC::LANDRY 21 lines 28-JUL-1995 18:54
-< Robot problem solved after a tough fight! >-
--------------------------------------------------------------------------------
Today, the system had a major problem or two..............
First off, the computer locked up. So, it was rebooted and
initialized. All looked well. Tried running a wafer through
the system, but the system came up with a robot malfunction.
The open sensor wasn't being made. Found out that the robot
was interfered with when trying to prevent a wafer mishap.
No problem, I'll just reset up the flags/robot positions.
Nay, nay! Not so easy. I had a heck of a time setting these
states up. After some long, grueling hours at it, all finally
fell into place. Ran several test runs successfully.
Second problem occurred as the robot fix was concluding. We
took an electrical hit from a storm, knocking the systems out.
Along with it went the DI water going off. The system was
rebooted immediately, but we had to wait 10-15 minutes for the
water to come back on.
Doc
|
45.42 | robot finger position error | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:10 | 15 |
| SUBPAC::LANDRY 12 lines 29-JUL-1995 18:03
-< robot finger position error >-
--------------------------------------------------------------------------------
The robot finger position errored today. See a couple of notes
back for Fridays Robot Finger Positioning happenings. Anyway, the
systyem cvouldn't recover due to the robot not recognizing it's
position. I attempted to manually move the robot around to recover the
wafer, but it didn't go well. The wafer broke in the spin station.
I cleaned up the five or six pieces it broke into (easy cleran) and
then rtebootyed the system. Ther robot worked well. Assured the the
flags were being made, which they were.
Doc
|
45.43 | PROCESS ENGINEERING ADJUSTMENTS CHANGED | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:10 | 13 |
| SUBPAC::LANDRY 10 lines 4-AUG-1995 19:24
-< major adjustments made per / Pro Eng >-
--------------------------------------------------------------------------------
After doing several expirements on A2 this morning, Frank Krupa asked
that the megasonics arm height be adjusted to ~3/8", along with the
adjusting of the generator to the max setting. We also adjusted the
megasonics flow rate meter to 500 ML (a setting of 5 on the flow
meter). These changes will improve the particle removement perform-
ance dramatically. Other changes will follow as the operations spec
needs to be DCO'd first.
Roger / Doctor
|
45.44 | carriage problems | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:11 | 35 |
| ASDG::POIRIER 11 lines 16-AUG-1995 20:09
-< SPin Station Back Plate Bracket Screws Stripped >-
--------------------------------------------------------------------------------
* Spin Station alignment problem:
Found the (back plate drive bracket) screws loose. We tried to
tighten them but they were stripped. The two 8-32 screw holes were
drilled-out and retapped to 10-32. The assembly was then reinstalled
and adjusted. The plexiglass cover needs to be mounted and a
verification run started.
John.
<<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]FILMS_ONLY.NOTE;1 >>>
-< Films Only! >-
================================================================================
Note 65.130 CMP-SCRUB.A1 130 of 145
SUBPAC::BJUBINVILLE "NFD Fire - Rescue" 13 lines 16-AUG-1995 08:37
-< Contamination and carriage problems >-
--------------------------------------------------------------------------------
Contamination source suspected from CMP.A1 brought skid
marks to the top brush in both stations. Both brushes were
changed and the passed MQC's.
Later this morning, the spin station grippers were not
fully grabbing the wafers. It was discovered that the carriage
was tilted and had loose screws. Once they were tightened, the
tilt got worse. After about an hour a screw had come lodged inside
the hole and needed TLC to get it out. 1st shift has taken over
and are looking into the fix now.
Bob and BJ
|
45.45 | spin and robot alignments, brushes replaced | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:11 | 19 |
| SUBPAC::BJUBINVILLE "NFD Fire - Rescue" 16 lines 16-AUG-1995 23:21
-< spin and robot alignments, brushes replaced >-
--------------------------------------------------------------------------------
The problem with the spin station was discovered by 1st shift.
The 10/32 screws that were used needed to be shorter so 1/2" screws
were installed. The spin station and carriage alignments are now ok.
The spin lock positioning was out 90 degrees. The spin station shaft
needed to be loosened up and repositioned. The robot was now out of
alignment so I went through part of the procedure with Roger. The
robot now picks up the wafers ok.
The bottom brushes were replaced per ENG. All (8) lower bearings
were replaced as well due to stalling. Wafers are being run now for
VERIFY and brush break-in and we've expericenced no problems at all.
The next reply will list all new current machine offsets.
Roger and BJ
|
45.46 | SYSTEM OFFSETS AS OFF 8/16 | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:12 | 14 |
| SUBPAC::BJUBINVILLE "NFD Fire - Rescue" 11 lines 16-AUG-1995 23:22
-< new system offsets >-
--------------------------------------------------------------------------------
The current System Offset settings are:
Load Station______1310 Rotation @ Spin Station______871
Unload Station_____490 Rotation @ Receive Sta._____1813
Wrist @ Spin Sta___110 Balde______________________11500
Wrist @ Receive____240 Robot Fingers________________340
Spin Xport Delay____40
|
45.47 | PROBLEMS WITH BRUSH HEIGHTS, EXHAUST AND NO AMMONIA | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:13 | 27 |
| SUBPAC::LANDRY 24 lines 23-AUG-1995 19:18
-< some training / findings info >-
--------------------------------------------------------------------------------
WED - 08/23/95
Ron Simmons from Ontrak worked on CMP-SCRUB.A1 today. John Poirier, Doc,
and Brendan Finn, and myself were trained by Ron to perform the following:
1 brush change procedure: NOTE: NEED A BUCKET IN CMP ROOM TO SUBMERSE
THE BRUSH CORE IN TO MAKE INSTALLATION
EASIER.
2 brush parallelism: WERE OUT
3 brush height set-up: BOTH BRUSHES WERE SET TOO HIGH
In addition Ron performed a "once-over" on the tool and identified the
following issues; some of which were corrected:
- No NH4OH was flowing to SCRUB.A1. All flow regulators were closed.
even when they were opened, there was no supply and NO ALARMS!
Doc called Ashland chemical. Shortly after the call we heard the low
flow alarm go off. Soon after the NH4OH began to flow. Plan to
look into this further tomorrow to determine how long this was this
way and why we had no alarms.
- The exhaust for the spin station and brush stations are not set-up
correctly(their required exhaust tags are actually reversed).
Brush station should be ~5 scfm. Spin station should be ~50scfm.
Actual was ~16 scfm for both.
Frank, Ron, Doc, John, Brendan
|
45.48 | BRUSH ROLLER ARM NOT OPENING, BROKE | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:13 | 15 |
| SUBPAC::LANDRY 12 lines 2-SEP-1995 19:38
-< Broke it's arm.......ouch! >-
--------------------------------------------------------------------------------
The system wouldn't release the wafer from the Outside Brush station
today. The left roller arm was broken. The shaft would actuate, but
the arm stayed stationary. Replaced the arm tested several times.
Back up and running.
Just a reminder.........the bolts that hold the arms to the shaft
should never be tightened to capacity. Over tightening will only
stress the material. Proper installation will leave a small gap at
the slit on the locking flange. No gap......No good.
Doc
|
45.49 | LEAKS IN THE BRUSH STATION TUBS | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:14 | 14 |
| SUBPAC::BJUBINVILLE "NFD Fire - Rescue" 11 lines 8-SEP-1995 07:13
-< Resealing time >-
--------------------------------------------------------------------------------
Removed both outside brush station roller drive asmbly.
and re-siliconed the fitting. It will be ok to use around
11:30am or so. The brushes are in the DI sink cascading.
The one with the green strap through it is the TOP brush.
The other two shafts in the inside station were siliconed
on the outside as well.
D shift
|
45.50 | brush station roller arm broke | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:14 | 11 |
| PATE::GEE 8 lines 20-SEP-1995 00:58
-< Inside brush station roller arm >-
--------------------------------------------------------------------------------
Problem : A wafer went past the inside brush station and stopped
before the spin station.
We found the left inside brush station roller arm broken
at the shaft. Replaced the arm and ran verification wafers,
all set.
B shift
|
45.51 | Robot finger flags loose | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 08 1995 22:14 | 11 |
| SUBPAC::BJUBINVILLE "NFD Fire - Rescue" 8 lines 27-OCT-1995 06:51
-< Robot finger flags loose >-
--------------------------------------------------------------------------------
The system kept alarming for fingers not closed. Upon
removing the covers to the robot, we observed the flags
were loose and needed tightening. The entire robot open/close
and wafer pick-up needed to be re-taught.
D shift
|
45.61 | wafer shattered in send | LUDWIG::RIDLON | Eliminate the obvious | Tue Dec 12 1995 06:48 | 10 |
| Wafer shattered in send portion of machine (product). There was only
one wafer left in the cassette when it happened. Tried to duplicate
problem by cross slotting a wafer but once the wafer reached the wafer
sense the boat lifted off its sensor and the elevator returned to home.
The only thing we can figure for a reason is that the wafer was chipped
or had stress cracks prior. Unit was cleaned and a cassette of dummies
ran through with no problems. The subsequent particle check passed.
B shift
|
45.62 | Scrubber NH4oH OCP's Jumped-out | ASDG::POIRIER | | Fri Dec 15 1995 19:34 | 16 |
|
* NH40H OCP Failure:
-----------------
The OCP failed at lunch time today stopping the flow of NH40H.
Ashland cannot find the problem. They are currently working with
Systems Chemistry to find a solution. In the mean time, we have
continued running both scrubbers with the OCP connection in the
OnTrak electrical cabinet JUMPED-OUT. The OnTraks cannot detect
any drop in NH40H under the current circumstances. Therefore,
the NH40H pressure must be visually checked before each production
run.
Please show all operators on your respective shifts the location of
the Nh40H gage and instruct them how to read it.
|
45.63 | fluke | LUDWIG::RIDLON | Eliminate the obvious | Sun Dec 17 1995 03:46 | 14 |
| System alarmed for robot fingers not closing. Wafer was there fingers
were closed but the flag wasn't being satisfied. The fingers were loose
so I opened and closed them a couple of times but still it wouldn't see
the closed flag. Went in to teach mode and increased the robot fingers
from 190 to 200 which closed them tighter and satisfied the sensor.
Opened and closed the fingers and everything was working but the
fingers seemed to be closing too tightly on to the wafer. Put the robot
fingers number back to 190. Opened and closed them and the play was
gone that was noticed before I changed the numbers in the first place
and the closed sensor was being satisfied. Tried it a couple of more
times in manual with the same results. Resumed operation and it
finished the lot with out alarming.
B shift
|
45.64 | feedback adjustment made to roller motor | SUBPAC::LANDRY | | Sun Dec 17 1995 19:08 | 8 |
|
The system alarmed several times today for "inside brush station roller
motor out of spec". The system never stopped running, just alarmed.
The rotation speed was checked with the tachometer. It was spinning at
41 RPM but the readout said 32-34. So adjusted the board potentiometer
for the readout to match the actual. All is well!
Doctor
|
45.65 | Inside Brush Station rollers | FABSIX::R_GEE | | Tue Dec 19 1995 07:03 | 16 |
| * Inside brush station rollers not turning.
- Found DI water dripping down onto the motor shaft.
Replaced the o-ring seal on the roller shaft.
Installed an o-ring on top of the teflon o-ring
housing to prevent the assembly from working up.
Replaced the motor.
Replace the motor driver board, it failed also.
Calibrated the spin speed, allset.
B shift
|
45.66 | SYSTEM OFFSETS AS OF 12/21/95 | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Dec 21 1995 00:36 | 19 |
|
LOAD STATION -- 1310
WRIST AT SPIN STATION -- 124
WRIST AT RECEIVE -- 240
SPIN XPORT DELAY -- 35
ROTATION AT SPIN STATION -- 873
ROTATION AT RECEIVE -- 1813
ROBOT FINGERS -- 190
UNLOAD STATION -- 490
BLADE LIFTER -- 11250
|
45.67 | ROBOT ROTATION ADJUSTED | FABSIX::B_WESTPHAL | | Fri Dec 22 1995 17:10 | 16 |
|
CMP-SCRUB.A1 - DOWN FOR NOT PICKING THE WAFERS RIGHT FROM THE SPIN STATION.
FOUND THE ROBOT TO BE GRABBING THE WAFERS TO HIGH ON THE FINGERS. THIS
PORTION OF THE FINGER IS ALITTLE THICKER, PREVENTING THE ROBOT FROM
CLOSING PROPERLY GIVING A "FINGERS NOT CLOSED" ERROR. ADJUSTED THE
ROTATION FROM 873 TO 870, REBOOTED THE SYSTEM RAN A BOAT OF WAFERS W/
NO FURTHER PROBLEMS. TOOL WAS RETURNED TO PRODUCTION. THE OFFSETS
ARE NOW AS FOLLOWS:
ROT. AT SPIN 870
ROT AT RCV 1813
WRIST AT SPIN 124
WRIST AT RCV 240
ROBOT FINGER 190
LIFTER 11250
UNLOAD STATION 490
LOAD STATION 1310
|
45.68 | SYSTEM OFFSETS AS OF 12/22/95 | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 22 1995 20:42 | 19 |
|
LOAD STATION -- 1310
WRIST AT SPIN STATION -- 124
WRIST AT RECEIVE -- 240
SPIN XPORT DELAY -- 35
ROTATION AT SPIN STATION -- 870
ROTATION AT RECEIVE -- 1813
ROBOT FINGERS -- 190
UNLOAD STATION -- 490
BLADE LIFTER -- 11250
|
45.69 | Start-up for 12/29 | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Dec 29 1995 15:11 | 8 |
|
The DI came on but is still out of spec so we didn't want
to keep it to the tool so we shut off the DI feed and opened the
by-pass. Facilities still have some by-passes opened, that's
why we don't have very high pressure.
Mitch,John,Leon,BJ
|
45.70 | CMP-SCRUB.A1 START-UP COMPLETE | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Sun Dec 31 1995 14:39 | 25 |
|
CMP-SCRUB.A1 STARTUP CHECKSHEET
-------------------------------
Shift Date
------- -------
1. Ensure that all necessary facilities are B/C/D 12/31/95
available and ready to use.
2. Power up each system. B/C/D 12/31/95
3. Open incoming DI Water Valve to each B/C/D 12/31/95
system.
4. Close UDR valve at each system. B/C/D 12/31/95
5. Replaced all PVA scrub brushes. B/C/D 12/31/95
6. Change RTC status to SCHEDULED DOWN and
perform a Weekly PM on each system.
7. Change RTC status to VERIFY and notify
Operations that the systems are ready
for MQC.
|
45.71 | Ammonia | FABSIX::R_GEE | | Tue Jan 02 1996 07:48 | 3 |
| * Ammonia was down from 8 pm to 3 am.
B shift
|
45.72 | multiple offset adjustments made | SUBPAC::LANDRY | | Wed Jan 03 1996 17:51 | 10 |
| The system faulted out this afternoon for "robot finger not closed".
Upon checking out the problem, the fingers were overdriving closed.
Began to adjust the offsets, when I noticed that the "robot at spin"
position wasn't making it to the wafer. So, adjusted the offsets for
that, but found the "wrist at spin" position off. Adjusted that
offset, readjusted the fingers closed offset and tested system
performance. Voila! Alls hummin' along well.
Doc
|
45.73 | Arm replaced | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Wed Jan 10 1996 06:20 | 4 |
|
Replaced rear top roller arm in the outside brush station.
BJ
|
45.74 | robot arm flag adjustment | SUBPAC::LANDRY | | Wed Jan 17 1996 16:35 | 14 |
| First thing this morning, the robot was failing to pickup the wafer
from the spin station. In fact, the fingers closed position was was
a good inch away from wafer contact. Upon check out of the system,
found that the open sensor was the actual culprit. I hate to say it,
but it appears that someone may have physically forced the arms open
at some point. Both flags were securely in-place, but the open flag
was 45 degrees out of position. Hhmmmm.....
It took several attempts to set it up to where both flags would make
during a normal cycle. Then the robot closed position was re-taught to
fine tune.
A Shift
|
45.75 | broken "exercise" wafer in spin station | SUBPAC::LANDRY | | Sun Jan 21 1996 19:07 | 14 |
| The system broke a wafer (of the exercise type, NOT product) in the
Spin Station today. It's unclear as to how it happened, as we don't
normally monitor each wafer through it. It most certainly wasn't
the robot due to the fact that the spin station lock had not yet
energized to secure the fingers in-place. It appears that the spin
fingers didn't secure the wafer properly, and that the high speed (or
dry cycle to some) sent it flying to it's death.
Anyway, the debris was removed, then vacuum'd, then wiped....rinsed..
wiped and finally rinsed once more. A cassette of wafers was run to
verify system performance. An MQC check was then performed and it
passed with excellent results.
A shift
|
45.76 | Door latch | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Feb 09 1996 05:16 | 6 |
|
Tool kept alarming because the door latch alignment was out and
couldn't latch. Repositioned latch and monitored a boat of wafers
without any problems.
D shift
|
45.77 | OCP failure | SUBPAC::LANDRY | | Tue Feb 13 1996 19:34 | 9 |
| Today, the OCP for the Ammonium Hydroxide kicked off, sending both
scrubbers into a halt mode. Checked the pressure gauge and it had
40 PSI going to the machine. So why did they alarm. Well, the
fault displayed a "timed out" (or something wordy like that). We
weren't able to reset the OCP's. Gave Ashland a call to check out
there end of the deal. The computer that controls the OCP hichup'd.
They need to down load a couple of files and we were back in business.
Doc
|
45.78 | 2 X slotting??? | SUBPAC::LANDRY | | Wed Feb 14 1996 19:26 | 10 |
| Late in the shift, a wafer broke in the spin station. The mess was
cleaned up, exercise wafers were run and MQC's were performed. The
system is up.
Now, as to why the wafer broke is a bit confusing. I was shown a
cassette with double slotted wafers. Not sure if was ingoing or
outgoing cassette. Anyway, exercise wafers ran like a champ.
Doc
|
45.79 | roller drive shaft | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Feb 15 1996 06:27 | 9 |
|
Water once again has plaqued our shafts. This time it
was the roller drive shaft in the inside brush station. DI
got by a defective o-ring and went down the shaft and freezing
up the bearings. The entire motor drive shaft needed to be
removed from the tool and cleaned. After everything was cleaned
up, the rollers ran smooth.
D shift
|
45.80 | Contact phone numbers for OnTrak | SUBPAC::TOOKER | | Fri Feb 16 1996 10:50 | 27 |
| Here is the latest list of service contacts for OnTrak Systems. These con-
tacts can be used to get answers to technical questions, field service requests,
and parts.
1) Ron Simmons: Office - 508-643-2772
Pager - 508-504-7906 (this is your best bet)
If no response in one hour, contact area supervisor in Fishkill, NY.
2) Pat Paino (area supervisor), Fishkill, NY:
Pager - 800-963-5770
If no response in one hour, contact area supervisor in Austin, TX.
3) Stan Jaskinski (area supervisor in Austin, TX:
Office - 512-448-4494
Pager - 800-905-9275
4) Brad Clements (regional manager), Austin, TX: (best bet for nights)
Office - 512-448-4494
Pager - 800-796-7363 PIN 193-3639 (7 X 24 response)
|
45.81 | IB Upper Brush Motor blues..... | SUBPAC::LANDRY | | Sun Mar 10 1996 18:59 | 35 |
|
Through most of the shift today, the system would occasionally
alarm for an Inside Brush Station Upper Motor out of spec.
The brush would spin as low as 85 RPM at times, other times
it would spin within the 10% limits.
Through indepth trouble shooting, it has been determined that
the motor for that brush is the culprit to the problem. WE
calibrated the board for the motor driver and were unable to
attain the proper spin speed. The screen readout and the hand
held tachometer coincided as low rotation (~ 85-95 RPM). WE
tried swapping motor driver boards with the outside brush station.
The problem stayed with the IB station upper brush, AND the
possible bad board was working fine for the outside brush station.
WE also checked swapped bearings between the upper and low brushes
with no change in where the problem still resided. So, WE took
a closer look at the motor. By applying pressure to the electronic
leads on the motor itself, we were able to change the rotation
speed. Again, proof that the motor is beginning to wimp out. The
leads are currently (no pun intended) positioned such that the motor
is spinning at ~115-118 with the set point at 120 RPM.
Of course storage cabinet VSR-1 in the stockroom is down. We are
unable to attain the motor from the cabinet (there are 3 in there).
WE have placed a call to the local OnTrak contact (Ron Simmons)
at his home here in lovely Massachusetts to see if he may have
a 30.3 VDC motor to lend us until we can get our hands on one.
We hope to receive a return call from him. He has been supplied
my (the Doctor) beeper number as well as the fab phone number in
his voice mail.
By the way, an MQC check was performed to make sure that the slower
rotation speed wasn't causing any problems. MQC's were real good.
A Shift
|
45.82 | Inside brush station top brush motor. | FABSIX::R_GEE | | Mon Mar 11 1996 07:05 | 5 |
| The inside top brush station motor failed. Replaced the
motor with one found in J.T.'s cabinet, we couldn't get
one from stock. No errors after the motor was installed.
B shift
|
45.83 | Lower brushes hesitating | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Mar 14 1996 23:37 | 7 |
|
Noticed during weekly PM that the lower brushes in
both scrub stations were hesitating while rotating.
Replaced all (8) delrin glass ball bearing to fix.
BJ
|
45.84 | NH4OH drip manifold tests | SUBPAC::KRUPA | | Wed Mar 20 1996 17:43 | 60 |
| 03/20/96 FJK & JT & Ron Simmons OnTrak
Subject: CMP Scrubber NH4OH drip manifold tests
Summary:
Today we tested the drip manifolds to deliver NH4OH to CMP-SCRUB.A1 in
place of the current spray manifolds. The benefit of the drip manifold is
a reduction of ~6X of NH4OH used compared to the spray manifold. Today's
tests consisted primarily of performing MQCs and comparing to std particle
results. Particles looked good with the drip manifolds. The wafers will be
submitted for mobile ion contamination testing prior to releasing the drip
manifolds to production.
Experimental:
MQCs were performed on CMP.B1 and wafers scrubbed on CMP-SCRUB.A1 using
the standard SPRAY manifolds and recipe. The SPRAY manifolds were then
replaced with the DRIP manifolds. Scrub only MQCs were performed to ensure no
mechanical problems were induced during install. Then MQCs were performed
again on CMP.B1 and wafers scrubbed on CMP-SCRUB.A1 using the DRIP manifolds
and a modified recipe(10sec added to NH4OH step and 10sec removed from H2O
step). Finally the DRIP manifolds were replaced with the SPRAY manifolds
pending the outcome of mobile ion contamination tests to compare the two types
of delivery manifolds. The results of all tests are tabulated below.
CMP.B1+CMP- CMP.B1+CMP-
SCRUB.A1 SCRUB.A1
REMOVAL %STD REMOVAL %STD
HEAD PER MIN DEV PER MIN DEV
---- ------- ---- ------- ----
1 3495 12.1 3216 8.6
2 3503 7.8 3166 4.0
3 3226 6.4 2845 3.1
5 3232 9.3 2902 6.4
CMP.B1+CMP- CMP.B1+CMP-
SCRUB.A1 SCRUB.A1
LARGE TOTAL LARGE TOTAL
HEAD DEFECTS DEFECTS DEFECTS DEFECTS
---- ------- ------- ------- -------
1 124 557 162 553
2 103 357 89 250
3 96 299 197 852
5 105 389 167 730
CMP-SCRUB.A1 ONLY: LARGE DEFECTS = 1; TOTAL DEFECTS = 18
Note: The fab was evacuated after the spray manifold test and prior to the
drip manifold tests. The pad on the polisher was not kept wet during this
period(~1hr). This may account for the lower removal rate and slightly higher
defectivity on the DRIP manifold wafers. Also, the initial MQC was too center
fast, so we adjusted the ID from 97 to 100 and the OD from 133 to 130 prior to
the drip test and is the cause for the reduced uniformity. Finally the defect
data reported is POST only data not DELTA. The SCRUB only results are right
in-line w/where the tool normally runs.
|
45.85 | Wafer break | SUBPAC::LANDRY | | Wed Mar 27 1996 18:34 | 14 |
| The system broke a wafer this afternoon in the Send Station.
There wasn't any equipment problems that caused this. It was
an unfortunate bit of luck with the Opts people. The cassette
was mistakenly joggled, hence the cassette present switch state
changed which allowed the lead screw to begin resetting. As
fate would have it, a wafer was beginning to exit the cassette
at this time. Crash!!! The wafer was dirven into the upper
section of the entrance mouth of the Outside Brush Station.
Thouroughly vacuum'd up the wafer fragments from the Send
Station and Outside Brush Station. Rinsed said station several
times. Ran several test wafers. Ran MQC's. All is well.
A Shift and special guest Hawk
|
45.86 | CMP-SCRUB.A1 shutdown completed for facilities tie-in | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Thu Mar 28 1996 06:17 | 28 |
|
SHUTDOWN: 5:00 AM ON 3/28/96
STARTUP: 9:00 PM ON 3/28/96
CMP-SCRUB.A1 SHUTDOWN CHECKSHEET
--------------------------------
Tech. Date
------- -------
1. Change RTC status to SCHEDULED DOWN. LEON 3/28/96
2. Remove all PVA scrub brushes and submerge
in DI water. If possible, DO NOT remove
the brushes from their cores. LEON 3/28/96
3. Perform a complete wipe down and dry of
all wet surfaces. LEON 3/28/96
4. Close main incoming DI Water valve to each
system and open the DI Water return valve
(UDR). This will allow all incoming DI Water
to flow to UDR. LEON 3/28/96
5. Power down each system.
LEON 3/28/96
|
45.87 | water leak | FABSIX::L_SHKOLNIKOV | | Thu Mar 28 1996 07:02 | 8 |
|
Inside brush station rear arm. Water coming through the motor shaft.
Removed old motor and shaft and resealed water leak, and changed
o-ring on white cap. Roller motor wasn't replaced - non of them
in the stock room 6.
D-shift
|
45.88 | Inside roller assembly | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Fri Mar 29 1996 07:23 | 18 |
|
This tool never came up from the shut-down.
The roller arms in the inside brush station wouldn't
stay closed thus letting wafers roll right on
through. The roller themselves were kept stopping.
The entire roller motor shaft assemble is junk.
I took it apart and broken bearings and washers
were everywhere. The teflon bearing that connects
the shaft to the motor was junk. The main shaft
has all kinds of wear marks and needs to be replaced.
Not many of the parts were in stock and not enough
to get things working again. I will speak to Jim T.
about getting this stuff rushed in....
BJ
|
45.89 | The parts for cmp-scrub.a1 will be here Saturday, 3/30/96 | SUBPAC::TOOKER | | Fri Mar 29 1996 15:05 | 8 |
| CMP-SCRUB.A1 is currently down waiting for parts. The parts have been ordered
and are scheduled to arrive on Saturday, 3/30/96. I have requested That
Security receive the parts from Federal Express at the HLO3 Security desk.
John Poirier and Brad Westphal will be here to install the parts. Please
call me at home on Saturday if there are any questions or problems.
Thanks,
Jim T.
|
45.90 | "Inside Brush Station" Shaft/Arm Assy Replaced | ASDG::POIRIER | | Fri Mar 29 1996 19:21 | 8 |
|
INSIDE BRUSH STATION SHAFT/ARM ASSEMBLY:
---------------------------------------
The Inside Brush Station Shaft/Arm Assembly has been removed
from WL-Scrub.A1 in the Wet Lab and installed into CMP-Scrub.A1
in CMP. CMP-Scrub.A1 is currently running verification wafers.
We will WUP the system after another sucessful run of 25 wafers.
This latest run will be completed at 7:25 p.m.
|
45.91 | inside brush station / wl_scrub.a1 | FABSIX::R_GEE | | Sun Mar 31 1996 08:09 | 5 |
| Completed the inside top brush station roller motor and
shaft assembly alignments. Re-assembled WL_SCRUB.A1
and ran 25 wafers, all looks good.
B shift
|
45.92 | Arm replaced | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Wed Apr 03 1996 08:09 | 5 |
|
Replaced broken lower front roller arm in the
outside brush station.
D shift
|
45.93 | Shaft and o-rings replaced | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Tue Apr 09 1996 13:30 | 7 |
|
Replaced front roller drive shaft and o-ring seal
in the outside brush station. we also replaced the
o-ring seal in the rear roller assembly in the same
station.
A shift + EE
|
45.94 | Wafer sensor and PIA board replaced | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Tue Apr 09 1996 19:56 | 7 |
|
The inside brush station entry wafer sensor went
south. It took out the PIA board for cable #1 and #2.
Replaced PIA board and sensor. Ran wafers and everything
was ok.
A shift
|
45.95 | Spin Station / inside brush station rollers | FABSIX::R_GEE | | Wed Apr 10 1996 08:07 | 15 |
| The system went down for outside roller arms not opening.
Found the bracket under the station that moves the arm assembly
had loosened up, tightened the bracket.
The spin station fingers were missing the wafers intermittently.
Found the wafer was being positioned in the spin station off
center. Adjusted the sensor offset to 35 and adjusted the wafer
exit position out of the inside brush station more to the back
of the system. This needed to be done to compensate for the drive
belt being stretched and needs to be run on the smaller groove.
This makes the wafer exit out of the station higher on the back
side and shoot the wafer slightly to the front of the system.
The system will run fine like this until new belts can be ordered.
B shift
|
45.96 | wafers not exiting outer station | FABSIX::B_FINN | | Tue Apr 16 1996 20:07 | 9 |
|
Wafers intermittently not leaving outer clean station. Arms moving
out, but not enough.
Adjusted outer station left roller to match inner station left roller
position. Ran wafers, tested aok.
A shift
|
45.97 | Motor speed errors, loose gear | SUBPAC::BJUBINVILLE | NFD Fire - Rescue | Wed Apr 17 1996 05:57 | 10 |
|
Tool kept alarming for outside roller motor speed
out of spec and no turning as well. Found the rear
pulley that is mounted to shaft to have loosened up and
rode up the shaft. This caused the belt to bind up and
cause the motor to stall. Re-aligned pulley and tested
ok....
BJ
|
45.98 | Inside Brush station | FABSIX::R_GEE | | Mon Apr 22 1996 08:09 | 12 |
| The system went down for the inside brush station roller
speed out of spec.
Found a water leak was dripping onto the motor that
rotates the rollers, the motor was full of water. We
replaced the o-ring that is in the shaft cap and installed
a second o-ring under the cap. Replaced the motor,
repositioned the motor, motor bracket and drive assembly so if
there is ever another leak it won't take out the motor.
Also, there was a crack in the poly housing around the
shaft feedthru, we put silicone around the outside of it.
B shift
|
45.99 | scheduled contractor work performed | SUBPAC::LANDRY | | Tue Apr 23 1996 19:52 | 20 |
| Today...........
The DI water input to the room was shut down to perform
some scheduled work with the DI line AND the OCP's had
work done concerning the communications/electronics
package. This necessitated us to remove ALL the scrub
brushes so that they may be placed in water to keep wet.
All brushes were labeled and re-installed in their original
station/position.
Ashland performed some electronic/communication work on the
OCP's. They verified that the work done to assure proper
communication between the delivery system/OCP/scrubber.
Once the water came back on and the system were allowed to
be started up again, the brushes were installed, test wafers
were run through the system, and MQC checks were performed.
The system came back up with no problems.
A Shift / Jim T
|
45.100 | motor and board change | STRATA::SWENSON | | Fri May 03 1996 16:03 | 15 |
|
Problem 1:: Inside brush station top brush stopped rotating.
Fix 1:: Replaced top brush motor. Calibrated D.C. Motor P.C.B.
Problem 2:: Upon boot up lost Wafer Transport Motion.
Fix 2:: Attempted to calibrate Wafer Transport P.C.B. and
found R31 to be non responsive. Swapped with known
good transport board and found problem to move with
original transport P.C.B.
Installed with new P.C.B. from stock and recalibrated.
Doc and Hawk
|
45.101 | double loading problem | SUBPAC::LANDRY | | Tue May 07 1996 20:35 | 27 |
| Today..........
I was summoned to the system to investigate why the system
was going into an alarm state. The alarm was for "looking for
Inside Brush Wafer" or something close to that. Observed the
system and found that the Outside Brush Station was releasing
the wafer, allowing the wafer to slide atop the wafer in the
Inside Brush Station. When I say release, I mean that the entire
station acted normally with the lifter, rollers, transport belt
all working like it was the normal time to release the wafer.
All this despite the IB station showing it's status as "IN PROCESS".
This happened 3 or 4 times within an hours running. This made
absolutely no sense what-so-ever. So, I checked all the boards
and their respective connectors for any loose/crooked connections.
I didn't find anything abnormal, although I did remove and reseat
several connectors.
Upon reboot, ran 60 wafers flawlessly. Had to give it back to them
as I couldn't reproduce the problem.
ps. Yes, I did try rebooting the system several times before looking
at the electronics. And yes, I did talk to the rep about the
outcome. He's never experienced this problem, but felt the
reseating was probably the fix.
Doc
|
45.102 | Outside brrr station. | FABSIX::L_SHKOLNIKOV | | Thu May 09 1996 07:43 | 9 |
|
*** Outside brush station - water leaks through the drive shaft
of the rear arm. Bearing parts was found under the shaft
assembly. Entire shaft assembly needs to be replaced.
We do not have replacement parts in stock. I will speak to Jim T.
about getting this parts rushed in....
Leon.
|
45.103 | Motors offset | STRATA::RIDLON | Eliminate the obvious | Mon May 13 1996 07:50 | 6 |
| The wafer rotation motors on the brush stations were relocated. Orings
are being used in the interim until the belts, which have been ordered,
arrive. The belt is the same used to raise and lower the indexer.
B shift
|
45.104 | The new NH4OH drip manifolds have been installed | SUBPAC::TOOKER | | Tue May 14 1996 17:16 | 7 |
| Today the new NH4OH drip manifolds were installed on the inside and outside
brush stations. The new flow meters have not been installed. If flow
adjustments are needed, adjust the flow meter until the NH4OH goes from
a fast drip to a steady stream. After the manifolds were installed, monitors
were cycled to verify operation and MQC's were performed.
Jim
|
45.106 | New NH4OH drip manifolds have been installed | SUBPAC::TOOKER | | Tue May 14 1996 17:18 | 7 |
| Today the new NH4OH drip manifolds were installed on the inside and outside
brush stations. The new flow meters have not been installed. If flow
adjustments are needed, adjust the flow meter until the NH4OH goes from
a fast drip to a steady stream. After the manifolds were installed, monitors
were cycled to verify operation and MQC's were performed.
Jim
|
45.105 | broken motor coupling | STRATA::FINN | | Tue May 14 1996 21:21 | 7 |
|
System was put down at 7:20pm for outside lower brush not rotating.
Found the 3 prong motor coupling nipple broken off. Brian is attempting
to locate or borrow one.
A shift
|
45.107 | Coupling broke/replaced | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Wed May 15 1996 08:26 | 7 |
|
The drive coupling for the lower brush in the outside
brush station was found broke on A shift. Leon pulled
the part from stock and replaced it.
BJ
|
45.108 | Inside brush station roller speed | FABSIX::R_GEE | | Mon May 20 1996 07:37 | 5 |
| The system errored for "Inside brush station roller speed
out of spec." With a setting of 40 rpm it was only rotating
at 32 rpm. We tached the roller speed and adjusted the dc
motor control board.
B shift
|
45.109 | Spin Station Slide Assembly Air Cyl Replaced | ASDG::POIRIER | | Fri May 24 1996 20:20 | 10 |
|
Spin Station Assembly hesitating during Z Axis travel:
-----------------------------------------------------
The Spin Station Assembly was moving in a slow "jerking" fashion
as it moved from the Middle to the Bottom position.
We replaced the 3 position Air Cylinder on the Spin Station
Slide Assembly. A 25 wafer verification run was performed with
no further problems.
|
45.110 | No top rinse | STRATA::RIDLON | Eliminate the obvious | Mon May 27 1996 07:29 | 4 |
| Noticed that the spin station top rinse wasn't coming on during
process. Isolated to faulty sv26. Replaced and tested fine.
B shift
|
45.111 | New brush heights / cylinder lift | SUBPAC::LANDRY | | Tue May 28 1996 20:28 | 23 |
|
Today, we started the system up from a shutdown situation. At this
time, new brushes were installed. The wafer count was zero'd out
in conjunction with the new brush install. MQC's passed on the 1st
pass.
The New Brush Heights are: IB-front 257 IB-rear 328
OB-front 315 OB-rear 350
The new brush heights have been entered into the latest Weekly PM
checksheet. The Weekly PM has been performed for this week.
The system is currently experiencing some difficulties with the Spin
Station lift cylinder. It occassionally alarms looking for the top
position sensor. Whats happening is that the spindle is remaining
at the bottom of the spin station, then comes up after a short pause.
This is inconsistant in that it happened 4-5 times over the course of
the day. It will need to be looked at.
A shift
A shift
|
45.112 | Spin station lift assembly | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Wed May 29 1996 07:33 | 15 |
|
Came into the shift with problems with the spindle
not always arriving to the top position and was alarming.
It was intermitting all morning but since there was tons
of work in the area, it was tolerable. Around 3am it was
failing after every wafer. I increased the top solenoid
air and backed off on the bottom solenoid air. After 50
wafers and no problems, it was turned of to OPS.
During this fix, a wafer had slid by the outside brush station
rollers. The front top roller arm was broke and needed to
replaced. That was taken care of.
BJ
|
45.113 | Outside Brush Station Motor & Bearings Replaced | ASDG::POIRIER | | Thu May 30 1996 20:41 | 11 |
|
Brush Bearings:
--------------
The Outside Brush Station Brush bearings were replaced on both the
top and bottom brushes.
Brush Station Motor:
-------------------
The Outside Brush Station (Lower Brush) Motor was replaced and
calibrated this afternoon.
|
45.114 | I.B. drive shaft replacement. | FABSIX::L_SHKOLNIKOV | | Sat Jun 15 1996 07:25 | 9 |
| INSIDE BRUSH STATION SHAFT assembly
------------------------------------
The Inside Brush Station Front Drive Shaft Assembly has been
removed from the tool and replaced with a new one. I also replaced
the o-ring seal inside of the white cap and checked roller
alignments. Scrub. A1 is currently running verification
wafers.
Leon
|
45.115 | carriage level adjustment | SUBPAC::LANDRY | | Mon Jun 17 1996 20:54 | 8 |
| The system alarmed a couple of times today for "lost or broken wafer in
the spin station". The spin fingers were missing the wafer at the back
grippers.....occasionally. Found the carriage to be angled slightly to
the front, hence angling the wafer to cause this mishap to happen.
Made a slight adjustment to level the carriage. No failures
re-occurred.
A
|
45.116 | heat lamp replaced....... | SUBPAC::LANDRY | | Tue Jun 18 1996 19:41 | 16 |
|
The Spin Station heat lamp burn't out today. Thye filament was
shattered into many pieces, not just a filament break. Ordered
a lamp from the stockroom, which showed an inventory of two spares
lamps. When attempting to pick up the lamp, the stockroom folks
said that there were no lamps available. Blah blah blah, blah blah.
So, they said they would rush in a lamp, due to arrive tomorrow
morning. In the mean time, we were granted permission to "procure"
the lamp that belongs to A3, which isn't hooked up yet. It was
installed, but a test run failed. The lamp wasn't firing up. The
problem was traced back to the 10A fuse, located in the electronics
cabinet. Replaced the blown fuse and re-tested. All is well.
When the new lamp comes in, it needs to be placed on A3.
A
|
45.117 | Brush station rollers | FABSIX::R_GEE | | Wed Jun 19 1996 04:56 | 9 |
| The system was intermittently double slotting wafers at the
inside brush station.
Adjusted the outside brush station rollers up, so the wafer
hits them in the center.
Adjusted both the inside and outside brush station roller arms
in, so the wafers stop centered.
MQC's passed.
B shift
|
45.118 | Roller shaft replaced | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Sat Jun 22 1996 08:14 | 10 |
|
INSIDE BRUSH STATION SHAFT assembly
------------------------------------
The Inside Brush station rear drive shaft assembly has been
removed from the tool and replaced with a new one. I also replaced
the o-ring seal inside under the white cap and checked roller
alignments. Scrub.A1 is currently running verification
wafers.
Leon
|
45.119 | spin to top adjusted | LUDWIG::RIDLON | Eliminate the obvious | Sun Jun 23 1996 06:50 | 14 |
| Alarming for "Timeout waiting for spin station at top sensor." Found
that the spin shaft would eventually make it to the top. Adjusted the
top, middle, and bottom solenoids and was able to correct the problem.
NOTE: Just an fyi, the solenoids for the spin shaft are labeled for the
direction they send the shaft not to where they are connected to the
bimba, ie., SPIN TOP sends the sation up which means its, obviously
connected to the bottom of the shaft. Another Fyi, for the TOP and
MIDDLE solenoids the bottom adjustment screws adjust the release of the
air from the bimba while the top screws adjust the application of air to
the bimba. The BOTTOM solenoid is the opposite. Clear as mud...right?
B-shift
|
45.120 | spin shaft solenoids | LUDWIG::RIDLON | Eliminate the obvious | Sun Jun 23 1996 22:32 | 7 |
| I was wrong....about the spin shaft solenoids.
The solenoids (and if you know this already good. I'm basically putting
this in here for my sake) have two different outputs; A and B. The top
screw adjusts the amount of air going to port A and the bottom screw
adjusts the amount of air going to port B. The "SPIN TOP" and "SPIN
MIDDLE" use port A and "SPIN BOTTOM" uses port B.
|
45.121 | panel / gauges installed | SUBPAC::LANDRY | | Tue Jun 25 1996 20:41 | 17 |
|
While the room was shut down for a couple of hours today, we
took advantage of that time and installed two inline gauges.
One for the DI Water and the other for Ammonia. The gauges
are mounted above the drip pan to the rear of the machine.
It is necessary to remove the rear panel (on the Send Station
end) to see the gauges. These are currently only for convien-
ience as they tie into a bigger project, which is alarm warnings
at the touch screen, which we don't presently have the capability
of on the systems today. New software and boards will be installed
in the near future which will give us that capability.
So, the two new gauges have been plumbed in place and are reading
the pressures. Be aware that 2% ammonia is flowing to it's new
gauge. The lines were checked for leaks, none were found.
A / Jim T
|
45.122 | Roller shafts replaced | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Thu Jun 27 1996 07:32 | 13 |
|
The front and rear roller shaft assemblies
in the outside brush station were replaced along
with the o-rings at the white shroud.
Upon disassembly, the rear lower arm was found
cracked so it was replaced.
All of the roller shafts in CMP-SCRUB.A1 have been
replaced. CMP-SCRUB.A1 will commence next.
D shift
|
45.123 | particles | STRATA::WESTPHAL | | Fri Jun 28 1996 09:30 | 8 |
|
SCRUB.1 - Down for failing MQCs for high particles. All brushes were swapped
out as requested by engineering. The brush heights now set at
OB front 315 OB rear 350 IB front 313 IB rear 383. All stations were
cleaned and rinsed w/ DI water. The IB exit rinse was not operating
so an adjustment was made to the appropriate AO valve. Particles
came in at 7. The tool was returned to production.
|
45.124 | OnTrak after hours support | SUBPAC::TOOKER | | Fri Jun 28 1996 11:28 | 6 |
| Here is a telephone number that can be used for contacting OnTrak Systems
field service and technical support after hours. This number will connect
you to an answering service. They will forward your message to the people
who are on call. The number is: 1-888-ONTRAK1
Jim T.
|
45.125 | power up o.k. | LUDWIG::SWENSON | | Fri Jul 05 1996 20:39 | 2 |
| Tool powered up with no issues
waiting for di water.
|
45.126 | one flow meter installed, one to go | SUBPAC::LANDRY | | Sun Jul 07 1996 20:26 | 13 |
| While performing the Weekly PM today...............
The Ammonia flow meter was changed out (a planned change) in the OB
station. This flow meter replacement is part of the previously started
retro-fit of the ammonia drip manifolds. The IB station still needs to
be retro-fitted in the future. Time will be scheduled.
Also today, the Weekly PM was performed. The splash guards for the top
brushes (in both stations) were missing. They were found and
re-installed.
Doc
|
45.127 | spin station carriage | FABSIX::R_GEE | | Sun Jul 21 1996 08:09 | 8 |
| The system errored for "wafer not found in the spin station".
The dummy wafer never made it to the spin station, it was stopped
between the inside brush station and the spin fingers. It looks
like the carriage rollers never started. While trying to recover
the wafer it broke in the spin station.
Cleaned the spin station and ran 25 wafers, no problems noticed.
MQC's came out good.
B shift
|
45.128 | Spin Station up travel | FABSIX::R_GEE | | Tue Jul 23 1996 07:52 | 4 |
| The system kept erroring for "spindle not arriving up".
Adjusted the air solenoid for spindle top and middle.
B shift
|
45.129 | di leak found/repaired | SUBPAC::LANDRY | | Mon Jul 29 1996 19:56 | 10 |
| Whilst performing the Weekly PM today.........
A DI leak was repaired. The leak, which has been un-identifiable for
a few weeks, was located at the fitting on the OB station drip
regulator. The fitting was te-taped and re-installed. No more water
was forming in the drip pan. Hopefully, this puts an end to the leak
situation.
A shift
|
45.130 | Spin station lift errors | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Thu Aug 08 1996 07:52 | 10 |
|
The tool was down at shift change for the spin station
not coming up when told. Replaced "little giant" air
cylinder and ran 50 wafers without any problems. The
scrubber ran fine for the rest of the morning without
any problems.
Quarterly PM completed
BJ
|
45.131 | Ammonia leak | FABSIX::R_GEE | | Sun Aug 11 1996 07:39 | 5 |
| A fitting behind the inside brush station ammonia flow
regulator was leaking. Removed the regulator and re-teflon
taped, all set.
B shift
|
45.132 | Do not use the rubber boots on the carriers | SUBPAC::TOOKER | | Fri Aug 16 1996 15:59 | 12 |
| Yesterday during Strasbaugh's visit, we learned that they no longer
recommend using the rubber boots on the carriers. Their data now
shows that the boot is interfering with the free movement of the
gyro bearings and therefore impacting uniformity. The boot was
designed to keep slurry from getting into the gyro bearings. Be-
cause the carriers are disassembled and cleaned every few hundred
wafers, not using the boot should not cause a problem. I have
gone through the cmp and carrier build room and removed all of the
boots.
Thanks,
Jim Tooker
|
45.133 | IB station wafer rotation problem | SUBPAC::LANDRY | | Mon Aug 19 1996 20:30 | 8 |
| The system has an intermittant problem with wafer rotation in the
Inside Brush Station. Found a broken arm in the station. This was
replaced and the results tested. It appears to be better, but not
up to it's recommended performance. Occassionally, a wafer
"flinches" as it's attempting to rotate. If someone gets the
opportunity tonight, please see what you can do with it.
Doc
|
45.134 | CM-SCRUB.A1 SHUTDOWN CHECKSHEET | FABSIX::R_POIRIER | | Mon Sep 02 1996 02:52 | 20 |
|
CMP-SCRUB.A1/.A1/.A3/.C1-7 SHUTDOWN CHECKSHEET
----------------------------------------------
Tech. Date
------- -------
1. Close main incoming DI Water valve to each
system and open the DI Water return valve
(UDR). This will allow all incoming DI Water
to flow to UDR. ___B___ _9/1/96_
2. Remove all PVA scrub brushes and submerge
in DI water. Label each brush with the
tool name. ___B___ _9/1/96_
3. Perform a complete wipe down and dry of
all wet surfaces. ___B___ _9/1/96_
4. Power down each system. ___B___ _9/1/96_
5. Enter checksheet into notesfile. ___B___ _9/1/96_
|
45.135 | NO TOUCH SCREEN CAPABILITIES | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Thu Sep 19 1996 08:01 | 23 |
|
Tool down at shift change for no touch screen capabilities.
The touch screen wouldn't let us go to the touch screen
calibration target. Tried cycling the power several times
without any luck. Swapped the following out with C-7 tool:
video card and cables
interface w/cables
ribbon cable to RS232 bulkhead
monitor
no luck....
I spoke with Patrick from OnTrak office in New York
and he we went the right way and surprised it wasn't one
of them. As a last shot, he had us swap out the CPU.
still didn't work. He suspects corrupted software and/or
the video driver issues. He is sending a REP A.S.A.P. out
of his office and someone will be here ~10:30am with laptop
and software in hand. Since Jim is out, he will contacting
Tim. in the a.m. as well.
D shift
|
45.136 | IS Brush Station Roller Arm / Transport Belt Slipping | ASDG::POIRIER | | Thu Sep 19 1996 20:32 | 10 |
|
Roller Arm:
----------
The Inside Brush Station Roller Arm was cracked. The Arm was replaced
and adjusted.
The Inside Brush Station Transprt Belt appears to be slipping
intermittently. D Shift is picking up the Belt issue.
C shift.
|
45.137 | Brush stations went out. | FABSIX::L_SHKOLNIKOV | | Fri Sep 20 1996 08:34 | 15 |
|
Came into the shift with problems with the outside station lower
brush had some hesitation as it rotated in idle speed and it would
not send the wafers in inside station. To fix that we replaced all
brushes bearings and both drive belts. Works fine now.
Inside brush station. The wafers were not spinnig but wobbling.
Found that the guide rollers were out of alignment and the set
screws which hold all shaft assembly in place were not tightened.
WE tightened set screws and spent some time on rollers alignment.
Run 75 verification wafers looks good. No alarms so far. Returned
tool to production. Needs full MQC's.
D shift.
|
45.138 | Outside brush station problems | LUDWIG::WESTPHAL | | Sat Sep 21 1996 20:45 | 6 |
| SCRUB1 - Down for a loud Squeak in the OB, OB upper motor errors and the wafer
riding over the top of the rollers in the OB station. Found the
Squeak to be coming from the back roller took the shaft apart and
cleaned and adjusted belt tension. The roller heights and motor issue
still needs to be addressed. Ran out of time.
|
45.139 | motor/rollers | LUDWIG::RIDLON | Eliminate the obvious | Sun Sep 22 1996 06:05 | 11 |
| Replaced the motor and calibrated. Read D-shift's passdown about the
rollers. We found the same problem. Rollers misaligned (the OB
station's rollers weren't even aligned with each other) and screws not
tightened. So who's adjusting these things without putting something in
notes? Readjusted the rollers and tightened the screws (but not too
tight, don't want to bust those fragile plastic arms). Ran some wafers
through that looked good.
B-shift
|
45.140 | Update on pressure interlock upgrades | SUBPAC::TOOKER | | Wed Oct 02 1996 17:27 | 9 |
| Today OnTrak field service continued the pressure interlock upgrade work on
the scrubbers. The work is complete on cmp-scrub.a3 except for tying into the
NH4OH supply. The tie in will require all scrubbers to shutdown for 30 min.
The work on cmp-scrub.a1 is still ongoing. The electrical work is complete,
but still have the plumbing left to do. When OnTrak field service leaves today,
cmp-scrub.a1 will be available for Operations to run product. Work will
continue tomorrow.
Jim
|
45.141 | UPGRADE | STRATA::WESTPHAL | | Wed Oct 02 1996 20:56 | 6 |
| SCRUB.A1 - Was at scheduled down today for upgarades. Sensors are being
installed to monitor pressures and check for proper operation of the
heat lamp. The Ontrack Rep. will return tomorrow to install
software and work on .A2. Tool presently sets in an up status
needing only MQCs.
|
45.142 | Update on pressure interlock upgrades | SUBPAC::TOOKER | | Fri Oct 04 1996 17:03 | 15 |
| Today OnTrak field service continued work on the pressure interlock upgrades.
They have now completed all necessary work on cmp-scrub.a1/.a2/.a3 except
for the following open items:
1) CMP-SCRUB.A1: Needs new software installed to enable the new pressure
& interlocks. The new software will arrive on Monday and
CMP-SCRUB.A2 OnTrak will be back on Tuesday to install.
2) CMP-SCRUB.A3: Needs final connection of the NH4OH pressure interlock
to the incoming NH4OH supply line. We will do this
ourselves because all three scrubbers will need to go
down for 45 minutes to perform this task.
Jim
|
45.143 | broke a wafer | STRATA::RIDLON | Eliminate the oblivious | Tue Oct 08 1996 06:26 | 7 |
| Dropped and broke a product a wafer. Found that the robot was too high
over the lifter and the wafers were dropping into the blade. Increased
the "robot at rcv" calibration by 5 counts. Observed several wafers
unload and all looked good.
B-shift
|
45.144 | Alarms tied in / software changed | SUBPAC::LANDRY | | Tue Oct 08 1996 20:35 | 9 |
|
New software was installed today. The new software has allowed
for the DI Water and Ammonia to be tied into the alarm package.
Both fluids were tested to assure that the alarms worked when
a lack of fluid pressure would occur. The software also ties
in the Heat lamp. If the lamp fails to illuminate, the alarm
sounds. This featiures was tested as well.
A shift
|
45.145 | Ignore this alarm | LUDWIG::WESTPHAL | | Thu Oct 10 1996 20:34 | 6 |
|
SCRUB.A1 - Both scrubbers have been alarming been alarming intermittently for a
"leak detected in machine floor". The rep believes the problem is
being caused by a floating ground. This does not effect the proper
operation of the tool. If the error is received just clear and
continue running. ONTRAK is working on a fix.
|
45.146 | OCP alarm | FABSIX::B_FINN | | Tue Oct 15 1996 15:41 | 6 |
|
OCP alarm for NH4 no flow. Flow was good, tried to reset OCP,
but could not, had to emo and reboot OCP controller
A shift
|
45.147 | LOWER DI OUT OF SPEC ALARM | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Thu Oct 17 1996 00:02 | 10 |
|
Tool keeps alarms for "low DI lower flow value."
Incoming DI at 40psi and all DI manifolds flowing fine.
Logged a WARNING on tool to ignore it at this time until
E.E. can address a possible alarm setting tolerance that
might need to be adjusted.
BJ
|
45.148 | DI Water pressure switch setpoint lowered | SUBPAC::TOOKER | | Thu Oct 17 1996 15:56 | 8 |
| To try to minimize the number of false alarms for low DI Water pressure we are currently seeing on cmp-scrub.a1, I
lowered the pressure switch setpoint to 15 psi. The new software that Ontrak is working on will give us a time
delay before there is an alarm. This time delay will allow for the normal pressure fluctuations that happen when
the solenoid valves are opening and closing without the false alarms for low pressure. I will be removing the
warning from this tool.
Thanks,
Jim T.
|
45.149 | The expanded version.. :^) | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Thu Oct 17 1996 20:11 | 18 |
| <<< ASDG::DISK$SYS_DATA:[NOTES$LIBRARY]CMP.NOTE;1 >>>
-< CMP >-
================================================================================
Note 36.148 CMP-SCRUB.A1 EE/PE PASSDOWNS 148 of 148
SUBPAC::TOOKER 8 lines 17-OCT-1996 14:56
-< DI Water pressure switch setpoint lowered >-
--------------------------------------------------------------------------------
To try to minimize the number of false alarms for low DI Water pressure we
are currently seeing on cmp-scrub.a1, I lowered the pressure switch
setpoint to 15 psi. The new software that Ontrak is working on will give us
a time delay before there is an alarm. This time delay will allow for the
normal pressure fluctuations that happen when the solenoid valves are
opening and closing without the false alarms for low pressure. I will be
removing the warning from this tool.
Thanks,
Jim T.
|
45.150 | Timeout Errors - Outside Brush Station | ASDG::POIRIER | | Fri Oct 18 1996 20:22 | 13 |
|
* "Timeout at Outside Brush Station" Errors:
-----------------------------------------
The frequency of these errors has increased dramatically in the
last hour. It looks like the Send Station drive belt is not
moving the wafer into the Outside Brush Station every cycle.
A slight tilting of the cassette sends the wafer thru. All 3
scrubbers are busy moving product at this time. Please adjust
at the first available opportunity.
C Shift.
|
45.151 | BEARINGS SHOT IN SEND STATION | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Fri Oct 18 1996 22:49 | 14 |
|
Tool down at shift change for OUTSIDE BRUSH STATION TIMEOUT.
Both pulley assemblies in the SEND station needed to have most
of the bearings replaced and one shaft replaced. The shaft
closest to the exit of the SEND station was and still is slightly
bent. There was none in stock. One will have to be ordered.
The teflon bracket on the tower closest to the exit has a small
hairline crack at the "fork" end at the front roller. This will
eventually need to be replaced as well. Tool ran a boat of wafers
with no problems. Tool went over to OPS for MQC's.
D shift
|
45.152 | new brushes / height settings | SUBPAC::LANDRY | | Mon Oct 21 1996 19:08 | 11 |
| Hi all,
Just a quick note to let you know that the brush height
readings for cmp-scrub.a1 have changed. The brushes were replaced
this week per process engineerings request, hence the necessity to
change the micrometer settings. The new settings are:
IB - F = 345 OB - F = 270
B = 345 B = 310
Thanks.... the A shifters
|
45.153 | Recieve Station Lifter Did Not Rise | ASDG::POIRIER | | Sat Nov 02 1996 19:23 | 11 |
|
Lifter failure:
--------------
The recieve station lifter failed to rise into position twice
this afternoon. The tool was powered down and rebooted. The
problem did not return.
C Shift.
|
45.154 | Semi Annual Incomplete Checksheet Posted. | ASDG::POIRIER | | Thu Nov 07 1996 19:38 | 5 |
| Semi Annual Incomplete Checksheet Posted. There is one item remaining
on the semi annual p.m. Clarification is needed on the Brush Rotation
RPM procedure.
C Shift.
|
45.155 | Some shorted wires. | FABSIX::L_SHKOLNIKOV | | Fri Nov 15 1996 23:32 | 10 |
|
There was a problem with the send station wafer sensor. Timeout
looking for cassette was going forever. Two out of three wire
connectors which located outside of send station were shorted out
to the metal pedestal of the sender and connector itself was
sitting in water. The connector pins were replaced with new.
Tool runs fine since then.
Leon.
|
45.156 | Low Ammonia alarm problem solved | SUBPAC::LANDRY | | Mon Nov 18 1996 19:19 | 7 |
| After a couple of days of intermittent "low ammonia" alarms, Ashland
was able to find the fault. They traced the problem to a stuck valve
within their control panel down stairs. The problem was fixed before
10:00 AM, with no alarms happening the rest of the shift. Hopefully,
it stays that way.
A shift
|
45.157 | low ammonia | FABSIX::M_RIDLON | Eliminate the oblivious | Tue Nov 19 1996 21:03 | 8 |
| Tool alarmed for low ammonia pressure. Gauge indicating a pressure of
around 20psi with occasional dips below the 18psi setpoint. Called
ashland to investigate and they couldn't find anything wrong but said
they did cycle between pumps which brought the pressure up to around 25
psi. Question whats a normal pressure we should be seeing?
B-shift
|
45.158 | low ammonia | FABSIX::M_RIDLON | Eliminate the oblivious | Wed Nov 20 1996 01:35 | 8 |
| Had a few instances of low ammonia pressure after Ashland left. The
pressure would go to 20psi and the dip below the setpoint momentarily.
If 25psi is the normal pressure perhaps the low setpoint should be
lowered to compensate for these momentary dips in pressure.
B-shift
|
45.159 | Regulator tie-in | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Fri Nov 22 1996 06:55 | 9 |
|
Tool is in BGN_EXTERNAL for contactors to install
a DI regulator. The brushes are in the white containment
unit on the back table by the sink. All (4) brushes are
clearly marked.
D shift
|
45.160 | Wafer sensor in spin station | SUBPAC::BJUBINVILLE | CMP-6 Equipment Engineering | Sat Nov 23 1996 01:35 | 7 |
|
The tool was alarming for "timeout waiting for wafer in
spin station." The wafer was there but the system wasn't
seeing it. Found the wafer sensor amplifier wasn't set
properly and needed to be tweaked.
BJ
|
45.161 | fiber myopia | FABSIX::M_RIDLON | Eliminate the oblivious | Wed Dec 04 1996 03:40 | 8 |
| Timeout waiting for wafer in spin station alarm again. Too many alarms
with too many tweeks, something else must be wrong. Checked the fiber
optic cable with a flash light and sure enough found that some of the
fibers weren't transmitting light. Replaced cable and recalibrated
amplifier.
B-shift
|
45.162 | marine motors required. | FABSIX::M_RIDLON | Eliminate the oblivious | Wed Dec 04 1996 03:53 | 6 |
| Inside brush lower roller alarming for out of spec low. Replaced the
bearings but it still alarmed. Pulled the motor off and it was full of
water. Replaced and recalibrated.
B-shift
|
45.163 | sonic doulbe trouble | FABSIX::M_RIDLON | Eliminate the oblivious | Mon Dec 09 1996 07:05 | 11 |
| Tool alarmed for Mega sonic low flow. Found sv28 not opening when
enabled. Replaced and flow worked fine, however, the arm would not
sweep correctly. In fact the sweep was quite erratic. Sometimes it
would sweep to the center and return BUT not to home and then other
times it would start to sweep and 5 seconds into it (arbitrary number)
and then the flow would turn off and the arm would stop. Found that the
PIA card for the Mega arm sweep function to be faulty. Replaced and
recalibrated the sweep home sensor and it worked fine.
B-shift
|
45.164 | Arm Home Error | FABSIX::J_SWENSON | | Thu Dec 19 1996 19:16 | 10 |
| Megasonic Arm Home Fault
- Increased the gain on the Megasonic arm home amplifier
- Ran 50 wafers, could not repeat error
- Tool ran the shift with no errors
C-Shift
|
45.165 | Low Ammonia Alarms / Shut Off Valve Stripped | ASDG::POIRIER | | Fri Jan 10 1997 19:17 | 25 |
|
Low Ammonia Alarms:
------------------
The system errored repeatedly for "Low Ammonia" today. The new gage
that is connected to the alarm was reading from 9 psi to 20 psi
during the alarm event. (The other two scrubbers were reading a steady
40 psi on their respective alarm gages.)
Ashland checked everything downstairs and it was AOK. We attempted to
replace the alarm gage but the teflon shut off valve was stripped.
It may be stuck half way open, we have no way of knowing. The knob
and shaft turns in the valve but the pressure does not change. The
shaft turns indefinitely in both directions. The nearest shut
off valve for the Ammonia is in the pipe gallery.
Jim Tooker wants the valve replaced during the next P.M. We reduced
the Ammonia flow on both brush stations from a steady stream to a
steady drip. This seems to have stopped the tool from repeatedly
alarming for Ammonia Flow. The tool has been placed on WARNING until
the shut off valve is replaced and the ammonia flow is increased.
C Shift.
|
45.166 | valve replaced | FABSIX::M_RIDLON | Eliminate the oblivious | Mon Jan 13 1997 07:26 | 5 |
| Replaced the ammonia shut off valve during the weekly. The one removed
was all the way open.
B-shift
|
45.167 | Meagsonics sweep errors | FABSIX::R_POIRIER | | Sun Jan 19 1997 07:08 | 11 |
|
We had a problem with the Meagsonics sweep last nite.
It would not home after a sweep. Tried adjusting the
gain and it worked for a little while. Then error would
come back. Found that the home sensor was to close to
the base plate. Adjusted sensor up off the plate and
sweep worked fine the rest of the nite.
B shift
|
45.168 | Status of new software rev. | SUBPAC::TOOKER | | Wed Jan 22 1997 15:39 | 8 |
| OnTrak field service came in today to install a new revision of
software. This new revision will add a small time delay before
alarming for low DI Water or NH4OH pressure. This will help to
eliminate alarms due to pressure fluctuations from valves actuating.
However after doing the installation, OnTrak found a bug in the
software. He removed the new rev. and re-installed the old rev.
Jim T.
|
45.169 | Possible cure to high particles | SUBPAC::LANDRY | | Tue Jan 28 1997 19:24 | 12 |
| Whilst checking the system today, it was noticed that the megasonics sweep
arm was not sweeping to it's proper dimensions. The actual sweep distance
was fine, as that motor is rotating fine. The air cylinder, which positions
the megasonics arm at the edge of the wafer prior to the sweep motor turning
on, was found to be mechanically wedged. This was caused by the backing out
of a threaded rod (which is used to make the connection with the pivoting
portion of the assembly). Re-threaded the rod into the block. Re-attached
pancake valve and associated bracketing. Tested and achieved positive results.
We'll monitor the system occassionally to assure that the rod stays in place.
A shift
|
45.170 | lower brush bearings installed | FABSIX::J_SADIN | Freedom isn't free. | Mon Mar 17 1997 19:34 | 9 |
|
Replaced outside-station lower brush bearings today due to lower
brush intermittently stalling. Running ok now.
a-shift
|
45.171 | Particles.... | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Fri Mar 21 1997 08:38 | 27 |
|
The tool seemed to be trending on the high side
for particles on the MQC's for the last two months.
First off, the megasonic unit was turned off.
All of the handling and alignments were checked and
seemed to be ok. MQC's were run and the tool is
putting on 50 adders. The bottom brush motor was
intermittently stalling and would sometimes stop.
It could be started again by simply touching the
motor. We pulled the motor and found all kinds of
carbon build-up within the drive shaft and coming
out of the motor as well. The motor and bearings
were replaced and the area cleaned up. There some
fine particles embedded in both brushes so were
ended up replacing all (4) brushes. After 50 wafers
were run, an MQC was run and it was slightly out
of spec so I started up another 50 wafers. The tool
started alarming for Timeout For Wafer In Outside Brush
Station. The transport assembly wasn't turning. The
main drive belt for that transport was replaced and
got things running again. We are running 50 wafers
now and we'll be running MQC's when they are done.
We'll see then if we found the problem. After 24 hrs.,
the brush heights will have to be re-checked and maybe
adjusted.
D shift
|
45.172 | particles | FABSIX::J_SWENSON | | Fri Mar 21 1997 19:53 | 36 |
|
PARTICLE PROBLEMS
-----------------
- tool was cleaned from send to recieve.
- amonia was adjusted to proper flow
- 100 wafers run through the tool
- particles failed @ 102
- replaced outside brushes
- ran 50 wafers
- particles failed @ 113
- adjusted brush heights
- left 310
- right 300
- ran particle check on scrub a3 @ 7
- ran last check on A1 particles @ 52
the particles are coming down another 50 wafers will need to be run
then the particles should come in.
will pass on to D- Shift
C-Shift
|
45.173 | Auto reply system never posted this from 3/22 | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Fri Mar 28 1997 07:22 | 34 |
|
Took over from C shift for the particles issue.
After 100 wafers were run, the tool still failed for
over 100 particles. All of the motors were pulled out
of the mounting shafts and there was carbon built-up on
all of them except the one motor we replaced yesterday.
There was signs of rust emitting from the outside brush
station transport assembly so that was all rebuilt with
new parts. Both roller arm shafts and many associated
worn parts were replaced. The inside station was in
fair condition as far as wear is concerned but no problems
were found. The megasonic arm at home needed to be
adjusted. When it came home, the arm was whacking up
against the front panel. The spray nozzle was bumped to
the left side of the wafer and wasn't spraying in the
middle of the wafer. 200 wafers were run through the
tool and MQC's were run... both wafers came in spec
and removed particles. pre-5-29-0 post-5-27-0
pre-6-32-0 post-5-24-0
The tool was turned over to OPS but please monitor the
MQC readings for a while to see the results stay near
zero. The brush heights ad been changed since C shift
so below are the new numbers.
D shift
OB IB
---- ----
The new brush heights are: Front 340 345
Rear 350 345
|
45.174 | Roller motor brush motor issues | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Fri Mar 28 1997 08:43 | 17 |
|
Tool initially went down for wafers not rotating
in the outside brush box. Upon discovery, the front
roller arm was out of alignment and is still being
worked on by C shift. While checking the alignments,
the bottom roller was rotating whenever it felt like
it. It was spinning at erratic speeds when set to
118 RPM's and was slowly rotating when told to stop.
A motor controller board was checked and we found the
jumpers in the wrong spots. The channel was set up a
a non-programmable control versus a programmable. The
jumper settings for that channel were suppose to be
J2, J3 and J8, not J1, J3 and J8. After the jumpers
were set and board cal'ed, the bottom brush worked
fine.
D shift
|
45.175 | o-rings (again) | FABSIX::J_SADIN | Freedom isn't free. | Tue Apr 01 1997 20:38 | 8 |
|
replaced outside brush station roller o-ring (send side) and inside
brush station roller o-ring (receive side). These were the large
o-rings that connect to the pulley on the transport roller motor
(wafers were not being inducted/sent).
a-shift
|
45.176 | Receiver/heat lamp | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Fri Apr 04 1997 07:00 | 12 |
|
Tool alarmed for "robot waiting for receive station."
The boat/receiver never RESET and manually resetting
wouldn't work either. The cassette present switch was
ok as well. Ok.. Ok.. get to the point... I powered
down the tool and I heard a "pop." Nothing seemed wrong
until I went to run some wafers and the tool alarmed for
no heat lamp. Guess what the pop was. The lamp filament
blew. After the lamp was replaced... we were good to go...
D shift
|
45.177 | Robot Dropped Wafer / Lifter Did Not Rise Up | ASDG::POIRIER | | Thu Apr 10 1997 20:28 | 11 |
|
Recieve Station Lifter:
----------------------
The robot dropped a wafer into the base of the tool due to the lifter
not rising up. This event happened around 6:00 p.m. The problem hasn't
reoccured since the reboot.
C Shift.
|
45.178 | Wafer dropped by robot | FABSIX::R_POIRIER | | Mon Apr 14 1997 08:41 | 13 |
|
Scrubber robot dropped a wafer after the
spin cycle. We were able to recover and
continue. This happened 2-3 times thru
the night. Due to problems with the other
scrubbers we didn't have the time to try
and figure out the root problem.
B Shift
|
45.179 | Needed a reboot | FABSIX::B_WESTPHAL | | Thu Apr 17 1997 20:28 | 4 |
|
SCRUB.A1 - Down for the robot not picking a wafer from the spin station. The
tool just needed the receive station reset.
|
45.180 | WAFER CRASHED | FABSIX::B_WESTPHAL | | Thu Apr 24 1997 21:11 | 8 |
|
SCRUB.A1- Down for smashing a wafer in the spin station. Cleaned up the crash
and wiped down the OB, spin, And unload stations. Noticed the
wafers were stopping in the spins station to early preventing the
spindle fingers from grabbing the wafer properly. Repositioned the
sensor and ran 20 wafers w/ no problems. Particles came in at 26 so
the tool was returned to production.
|
45.181 | Dummy wafer dropped in receive station | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Sat May 10 1997 09:13 | 6 |
|
Robot dropped a wafer in the receive station. Checked
lifter and robot handling and they were ok. Verified ok
with 25 wafers.
BJ
|
45.182 | robot teach functions changed | SUBPAC::LANDRY | | Sun May 11 1997 18:37 | 17 |
| We encountered some problems today with the robot not picking up the
wafer properly from the spin station. A wafer was placed in the
station, and we diagnostically checked out the functionalities. It
was discovered that there were several mis-aligned positions. First
off, the "robot position at spin" was off. The robot teach function
number for the "rotation at spin" setting was changed from 875 to 882.
This allowed the 3 robot grippers to align to all grab the wafer.
Further inspection showed that the grippers were grabbing the wafer,
but there was a difference between the back grip and the two front
grips of where the wafer was gripped. We then adjusted the "wrist at
spin" teach number to allow the grippers to be planar to the wafer.
Also, to fine tune the station, the robot gripper close teach number
was adjusted. It ran the entire day without further problems.
There is no "I" in........... A team
----
|
45.183 | Lifter issue.... | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Wed May 14 1997 07:30 | 33 |
|
First off - the tool still has higher particles during
MQC's. This is an on-going issue and that's why we've
been trying to get it.
Last night we noticed that certain product wafers don't
particularly like the outside brush station. The rollers
are trying to spin the wafers counter-clockwise (looking from
the front of the tool) but the brushes are forcing the wafers
to go the other way. This friction is causing wear on the
rollers and according the class I went to, this is a source
of particles. The workload was big this morning and CMP.A1
is going down for the monthly so I won't get to it.
Twice the tool dropped a wafer in the receive station.
The blade lifter never even came close to being high enough
to get the wafer. Therefore when the robot fingers opened
and tilted, there goes your wafer. I haven't seen any ASSIST
for this tool for dropped wafers but I know of at least three
occurrences within the last week between C and D shift alone.
After the second wafer I noticed the lifter wasn't high enough
and I tried STOPPING the tool but it was to late and the wafer
fell. It got a small scratch. I went into maintenance to
manually run the lifter but it wouldn't move. Then the screen
froze. I EMO'ed the tool but there was a wafer in the spin
station. The fingers had the wafer and the carriage was still
out so the wafer became two. I went through all of the lifter
assembly and I couldn't find anything wrong. 25 wafers were
run and there were no problems. The tool ran perfect for the
rest of the morning.
D shift
|
45.184 | robot/recievestation | FABSIX::J_SWENSON | | Thu May 15 1997 20:37 | 9 |
| Recieve station/robot issue.
tool locked up with wafer in spin station.
reset tool and could not recreate error so returned back to
production.
C-Shift
|
45.185 | Send station rebuilt | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Fri May 16 1997 06:25 | 7 |
|
Rebuilt entire send station with all new bearings,
belts ect. Found alot of rust, cooked bearings and drive
belts 50% stretched out than a new one. All alignments
were checked and verified. Tool passed MQC's at 23.
D shift
|
45.186 | DROPPED A WAFER | FABSIX::B_WESTPHAL | Bradley BP#661 EXT# 6610 | Sat May 17 1997 20:49 | 5 |
|
CMP-SCRUB.A1 - Dropped a wafer in the receive station. The station was reset
and 25 wafers ran w/ no problem so tool was returned to production
with an assist.
|
45.187 | MEGASONIC NOT AT HOME. | FABSIX::B_WESTPHAL | Bradley BP#661 EXT# 6610 | Wed May 28 1997 19:44 | 10 |
|
CMP-SCRUB.A1 - Down for "Megasonic arm not at home" errors and adjustments to
the receive station. While the tool was erroring The Megasonic arm
stayed over the spinning wafer waiting to see the arm sweep at home.
the sweep home sensor was adjusted and the tool ran fine. When the
problem first started I thought the receive cassette needed to be
reset. When I reset the station and the lifter blade missed the wafer
while checking for the wafer. This scratched it. The unload station
offset was adjusted to correct this situation.
|
45.188 | C shift passdown regarding particles | FABSIX::B_JUBINVILLE | | Wed May 28 1997 22:11 | 18 |
| From: ASDG::POIRIER "FAB6 CMP EQUIPMENT TECHNICIAN. PAGER # 254 FUNCTION 1. TEL:6610,5804,4774,6609." 28-MAY-1997 19:29:29.78
To: FABSIX::B_JUBINVILLE,FABSIX::L_SHKOLNIKOV
CC: POIRIER
Subj: passdown
Brian & Leon,
Here is the latest situation:
CMP-SCRUB.A1 - DOWN/UNSCHEDULED: Failed MQC'S for particles.
We were working in the spin station this afternoon and probably
contaminated the spin area. The PIA board for the megasonic sweep
motor was replaced.
C Shift.
|
45.189 | particles. | FABSIX::L_SHKOLNIKOV | | Thu May 29 1997 08:25 | 26 |
|
Came in with tool down for particles. After
running 50 wafers, MQC's were run and the tool
failed for 250 particles. Replaced brushes and
replaced the (4) smaller style rollers with the
original larger style roller. Ran a modified
brush break-in recipe and checked particles.
They came in at 115. Another 50 wafers were
run and the particles came in at 110. We
elimnated the megasonics and rinse and we
got the same results. We tried eliminating
each brush station and got pretty much the same
results but the outside station seemed to be
in question.
Here is the results we got:
spin only IB OB Stras.best
7-29-0 59-237-0 12-31-0 18-61-0 36-78-0
3-31-0 40-76 -0 23-61-0 26-88-0 30-91-0
We ran regular mqc and it failed again. 4-20-0 and after mqc 82-155-0.
D shift
|
45.190 | Particles | FABSIX::J_SWENSON | | Thu May 29 1997 20:31 | 28 |
| Particles
---------
- picked up from D-shift
- noticed spin station belt worn
- ran mqc wafer blocking spin staion belt
- came in spec with 33 adders
- replaced spinstation belt and wiped down entire
spin station upper and lower, also entire recieve
station.
- another mqc wafer and 0 adders were found
-------------
- towards the end of the shift had a lifter problem
which ops noticed.
- reset recieve station and ran 50 verification wafers
- no problems found, returned tool back to production.
C-Shift
|
45.191 | Bad boats..... | SUBPAC::BJUBINVILLE | CMP Equipment Engineering | Fri May 30 1997 07:56 | 19 |
|
Tool stopped with no errors. The boat in the
receive station was warped and the switch couldn't
be made. The boat was replaced and the wafer that
was in the boat was put into the new boat after
the station was reset. When the lifter was checking
to see if there was a wafer present, it missed the
1st wafer and dragged past it. The wafer seemed ok
and the PPS remarked the lot. The robot somehow
nailed the cassette and the finger flags got lost.
The robot sensors needed to be tweaked. The tool ran
25 wafers with out any problems and the tool was
brought back up.
D shift
This second boat was
found to be warped as well. Both hit the trash can.
|